### v1

# On the spectral norm of Rademacher matrices

Abstract.
We discuss two-sided non-asymptotic bounds for the mean spectral norm of nonhomogenous weighted Rademacher matrices. We show that the recently formulated conjecture holds up to $\log\log\log n$ factor for arbitrary $n\times n$ Rademacher matrices and the triple logarithm may be eliminated for matrices with $\{0,1\}$-coefficients.
1. Introduction and main results
One of the basic issues of the random matrix theory are bounds on the spectral norm (largest singular value) of various families of random matrices. This question is very well understood for classical ensembles of random matrices [2], when one may use methods based on the large degree of symetry. Recently, a substantial progress was attained in the understanding of unhomogenous models [13], especially in the Gaussian case [9, 3]. However, there are still many open questions in this area, the one concerning Rademacher matrices is discussed here.
In this paper we investigate the mean operator (spectral) norm of weighted Rademacher matrices, i.e., quantities of the form
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|:={\mathbb{E}}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j},$
where $(a_{i,j})$ is a deterministic matrix and $(\varepsilon_{i,j})_{i,j\geq 1}$ is the double indexed sequence of i.i.d. symmetric ±$\pm 1$ r.v’s.
Since operator norm is bigger than length of every column and row we get
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|\sim({\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|^{2})^{1/2}\geq\max\Bigl{\{}\max_{i}\|(a_{i,j})_{j}\|_{2},\max_{j}\|(a_{i,j})_{i}\|_{2}\Bigr{\}}.$
For two nonnegative functions $f$ and $g$ we write $f\gtrsim$ g (or $g\lesssim f$) if there exists an absolute constant $C$ such that $Cf\geq g$; the notation $f\sim g$ means that $f\gtrsim g$ and $g\gtrsim f$. Seginer [11] proved that for $n\geq 2$,
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\log^{1/4}n\Bigr{(}\max_{i}\|(a_{i,j})_{j}\|_{2}+\max_{j}\|(a_{i,j})_{i}\|_{2}\Bigl{)}$
and constructed an example showing that in general the constant $\log^{1/4}n$ cannot be improved.
In [8, Theorem 1.1] it was shown that for any matrix $(a_{ij})$,
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|$ $\gtrsim\max_{1\leq i\leq n}\|(a_{i,j})_{j}\|_{2}+\max_{1\leq j\leq n}\|(a_{i,j})_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}.$ (1)
Here and in the sequel $\mathrm{Log}\ x=\log(x\vee e)$ and $\|S\|_{p}=({\mathbb{E}}|S|^{p})^{1/p}$ denotes $L_{p}$-norm of a r.v. $S$.
It was also conjectured that bound (1) may be reversed, i.e., for any scalar matrix $(a_{i,j})_{i,j\leq n}$,
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|$ $\lesssim\max_{1\leq i\leq n}\|(a_{i,j})_{j}\|_{2}+\max_{1\leq j\leq n}\|(a_{i,j})_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}.$ (2)
The proof of [8, Remark 4.5], based on the permutation method from [9], shows that in order to establish (2) it is enough to show that for any submatrix $(b_{i,j})_{i,j\leq m}$ of $(a_{i,j})_{i,j\leq n}$ one has
${\mathbb{E}}\|(b_{i,j}\varepsilon_{i,j})_{i,j\leq m}\|\lesssim\max_{1\leq i\leq m}\|(b_{i,j})_{j}\|_{2}+\max_{1\leq j\leq m}\|(b_{i,j})_{i}\|_{2}+R_{B}(\mathrm{Log}\,m),$ (3)
where for a matrix $A=(a_{i,j})$ and $p\geq 1$ we put
$R_{A}(p):=\sup_{\|s\|_{2}\leq 1,\|t\|_{2}\leq 1}\Big{\|}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{\|}_{p}.$
Our first result states that this conjectured bounds holds for $\{0,1\}$-matrices.
Theorem 1.
Inequality (3) holds if $b_{i,j}\in\{0,1\}$ for any $i,j$. As a consequence, for any $E\subset[n]\times[n]$,
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|$ $\sim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}\mathbbm{1}_{E}(i,j)\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}.$ (4)
Inequality (3) for $\{0,1\}$-weights is a consequence of the more general Theorem 5 below, applied to the symmetric $2m\times 2m$0,1}-matrix $A=\left(\begin{array}[]{cc}0&A\\ A^{T}&0\end{array}\right)$. Estimate (2) follows from (3) as in the proof of [8, Remark 4.5].
Remark 2.
[8, Proposition 1.4] gives an equivalent (up to a constant) form of $R_{A}(p)$ for $\{0,1\}$-matrices:
$\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j}\mathbbm{1}_{E}(i,j)\varepsilon_{ij}s_{i}t_{j}\right\|_{p}\sim\max_{F\subset E,|F|\leq p}\|(\mathbbm{1}_{\{(i,j)\in F\}})\|.$
Hence the first part of Theorem 1 gives a positive answer to the question posed by Ramon van Handel (private communication):
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|$ $\lesssim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
$\phantom{aa}+\sup_{F\subset E,|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F\}})_{i,j}\|.$
One may also state the two-sided estimate (2) in the equivalent way as
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|$ $\sim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\max_{F\subset E,|F|\leq\mathrm{Log}\,k}\|(\mathbbm{1}_{\{(i,j)\in F,i,j\notin I\}})_{i,j}\|.$
Remark 3.
Two-sided bound on moments of norms of Rademacher vectors [7] gives that for every $p\geq 1$,
$\Bigl{(}{\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|^{p}\Bigr{)}^{1/p}\sim{\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|+R_{A}(p).$
Thus, the first part of Theorem 1 might be equivalently stated as
$({\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|^{2\lfloor\mathrm{Log}\ n\rfloor})^{1/2\lfloor\mathrm{Log}n\rfloor}$ $\sim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
$\phantom{aa}+\max_{F\subset E,|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F\}})_{i,j}\|.$ (5)
It is quite tempting to show (3) for symmetric sets $E$ via a combinatorial method, since for $n\times n$ symmetric matrix $A$ and $k=\lfloor\mathrm{Log}\,n\rfloor$, $\|A\|\sim(\mathrm{tr}(A^{2k}))^{1/{2k}}$. Such an approach worked for Gaussian matrices [4], but we were not able to apply it in the Rademacher case.
Remark 4.
Signed adjacency matrices were studied in [6] in connection with 2-lifts of graphs. [6, Lemma 3.1] shows that to each signed adjacency matrix of a graph $G$ one may associate the 2-lift of $G$ with the set of eigenvalues being the union of the eigenvalues of $G$ and of the signed matrix. Hence Theorem 1 provides an average uniform bound for new eigenvalues of random 2-lifts.
To state results for general matrices we need to introduce some additional notation. We associate to a symmetric matrix $(a_{i,j})_{i,j\leq n}$ a graph $G_{A}=([n],E_{A})$, where $(i,j)\in E_{A}$ iff $i\neq j$ and $a_{i,j}\neq 0$. By $d_{A}$ we denote the maximal degree of vertices in $G_{A}$. Observe that in the case of $\{0,1\}$-matrices $\sqrt{d_{A}}\|(a_{i,j})\|_{\infty}=\sqrt{d_{A}}\leq\max_{i}\|(a_{i,j})_{j}\|_{2}$.
Theorem 5.
For any symmetric matrix $(a_{i,j})_{i,j\leq n}$,
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}n)+d_{A}^{19/40}\|(a_{i,j})\|_{\infty}.$ (6)
Remark 6.
Since $\|(a_{i,i}\varepsilon_{i,i})\|=\max_{i}|a_{i,i}|$ we may only consider matrices with zero diagonal. Moreover, for any unit vectors $s,t$ we have
$\Big{|}\sum_{i\neq j}a_{i,j}\varepsilon_{ij}s_{i}t_{j}\Big{|}$ $\leq\|(a_{i,j})\|_{\infty}\sum_{i,j}\mathbbm{1}_{\{(i,j)\in E_{A}\}}\frac{1}{2}(s_{i}^{2}+t_{j}^{2})$
$=\frac{\|(a_{i,j})\|_{\infty}}{2}\Bigl{(}\sum_{i}s_{i}^{2}\sum_{j}\mathbbm{1}_{\{(i,j)\in E_{A}\}}+\sum_{j}t_{j}^{2}\sum_{i}\mathbbm{1}_{\{(i,j)\in E_{A}\}}\Bigl{)}$
$\leq d_{A}\|(a_{i,j})\|_{\infty}.$
Hence,
${\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{i\neq j\}}\varepsilon_{i,j})_{i,j}\|\leq d_{A}\|(a_{i,j})\|_{\infty}$ (7)
and it is enough to consider only the case $n\geq d_{A}\geq 3$.
The proof of Theorem 5 takes the most part of the paper. Bernoulli conjecture, formulated by Talagrand and proven in [5], states that to estimate a supremum of Bernoulli process one needs to decompose the index set into two parts and estimate supremum over the first part using the uniform bound and over the second part by the supremum of the Gaussian process. Unfortunately, there is no algorithmic method for making such a decomposition – a rule of thumb is that the uniform bound works well for large coefficients and Gaussian bound for small ones. We try to follow this informal recipe, decompose vectors $s,t\in B_{2}^{n}$ into almost "flat" parts and use the uniform bound when infinity norms of these parts differs. When they are of the same order we make some further technical adjustments (using properties of the graph $G_{A}$) and apply the Gaussian bound. The crucial tool in the estimating corresponding Gaussian process is an improvement of van Handel’s bound [12].
We postpone the details of the proof till the end of the paper and discuss now some of consequences of Theorem 5.
Theorem 7.
For any symmetric matrix $(a_{i,j})_{i,j\leq n}$,
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\leq\mathrm{Log}\mathrm{Log}(d_{A})\Bigr{(}\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n)\Bigl{)}.$
Proof.
Let $M:=\max_{i}\|(a_{i,j})_{j}\|_{2}$, $u_{0}=1$ and $u_{k}:=\exp(-(20/19)^{k})$ for $k=1,2,\ldots$. Let $k_{0}$ be the smallest integer such that $(\frac{20}{19})^{k_{0}}\geq\mathrm{Log}(d_{A})$. Then $k_{0}\sim\mathrm{Log}\mathrm{Log}(d_{A})$ and $u_{k_{0}}\leq d_{A}^{-1}$. We have
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|\leq{\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{|a_{i,j}|\leq u_{k_{0}}M\}}\varepsilon_{i,j})\|+\sum_{k=1}^{k_{0}}{\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{u_{k}M<|a_{i,j}|\leq u_{k-1}M\}}\varepsilon_{i,j})\|.$
For any $k$,
$d_{k}:=\max_{i}|\{j\colon\ |a_{i,j}|>u_{k}M\}|\leq u_{k}^{-2},$
so by Theorem 5
${\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{u_{k}M<|a_{i,j}|\leq u_{k-1}M\}}\varepsilon_{i,j})\|\lesssim M+R_{A}(\mathrm{Log}\,n)+d_{k}^{19/40}u_{k-1}M\lesssim M+R_{A}(\mathrm{Log}\,n).$
Moreover, using again Theorem 5
${\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{|a_{i,j}|\leq u_{k_{0}}M\}}\varepsilon_{i,j})\|\lesssim M+R_{A}(\mathrm{Log}\,n)+d_{A}^{19/40}u_{k_{0}}M\lesssim M+R_{A}(\mathrm{Log}\,n)\qed$
Remark 8.
In Theorems 5 and 7 we do not assume the symmetry of $(\varepsilon_{i,j})_{i,j}$. However analogous bounds holds for ${\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j})_{i,j}\|$, where $(\tilde{\varepsilon}_{i,j})_{i,j}$ is the symmetric Rademacher matrix (i.e., $\tilde{\varepsilon}_{i,j}=\tilde{\varepsilon}_{j,i}=\varepsilon_{i,j}$ for $i\geq j$), since
${\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j})_{i,j}\|$ $\leq{\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j}\mathbbm{1}_{\{i\leq j\}})_{i,j}\|+{\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j}\mathbbm{1}_{\{i>j\}})_{i,j}\|$
$={\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j}\mathbbm{1}_{\{i\leq j\}})_{i,j}\|+{\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j}\mathbbm{1}_{\{i>j\}})_{i,j}\|\leq 2{\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|.$
Obviously, $d_{A}\leq n$, so Theorem 7 (together with the standard symmetrization argument) implies that bounds (3) and (2) hold up double logarithms of $n$. However, decomposing matrix into two parts and using the Bandeira-van Handel bound one may derive conjectured upper bounds up to triple logarithms.
Theorem 9.
For any matrix $(a_{i,j})_{i,j\leq n}$,
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\mathrm{Log}\mathrm{Log}\mathrm{Log}\,n\Bigl{(}\max_{1\leq i\leq n}\|(a_{i,j})_{j}\|_{2}+\max_{1\leq j\leq n}\|(a_{i,j})_{i}\|_{2}+R_{A}(\mathrm{Log}\,n)\Bigr{)}$
and
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\mathrm{Log}\mathrm{Log}\mathrm{Log}\,n\Biggl{(}$ $\max_{1\leq i\leq n}\|(a_{i,j})_{j}\|_{2}+\max_{1\leq j\leq n}\|(a_{i,j})_{i}\|_{2}$
$+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}\Biggr{)}.$
Proof.
Assume first that the matrix $(a_{i,j})$ is symmetric. Let $g_{i,j}$ be iid $\mathcal{N}(0,1)$ r.v’s. The result of Bandeira and van Handel [4] implies
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim{\mathbb{E}}\|(a_{i,j}g_{i,j})_{i,j\leq n}\|\lesssim\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}+\sqrt{\log n}\|(a_{i,j})\|_{\infty}.$ (8)
Put $M:=\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}$. Estimate (8) yields
${\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{|a_{i,j}|\leq M\mathrm{Log}^{-1/2}\,n\}}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}.$
We have
$\max_{i}|\{j\colon\ |a_{i,j}|>M\mathrm{Log}^{-1/2}\,n\}|\leq\mathrm{Log}\,n,$
hence Theorem 7, applied to a matrix $(a_{i,j}\mathbbm{1}_{\{|a_{i,j}|>M\mathrm{Log}^{-1/2}\,n\}})_{i,j\leq n}$ implies
${\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{|a_{i,j}|>M\mathrm{Log}^{-1/2}\,n\}}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\mathrm{Log}\mathrm{Log}\mathrm{Log}\,n\Bigl{(}\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n)\Bigr{)}.$
Therefore, for any symmetric matrix $(a_{i,j})$,
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\mathrm{Log}\mathrm{Log}\mathrm{Log}\,n\Bigl{(}\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n)\Bigr{)}.$ (9)
Now, supppose that matrix $(a_{i,j})$ is arbitrary. Applying (9) to the symmetric $2n\times 2n$ matrix $\left(\begin{array}[]{cc}0&A\\ A^{T}&0\end{array}\right)$ we get the first part of the assertion.
The second part follows from the first one as in the proof of [8, Remark 4.5]. ∎
Organization of the paper. In Section 2 we discuss basic tools used in the sequel, including an improvement of the van Handel bound for norms of Gaussian matrices from [12]. In Section 3 we derive a weaker version of Theorem 7 with $\log(d_{A})$ instead of $\log\log(d_{A})$ factors. The last section is devoted to the proof of Theorem 5.
2. Tools
We will use the following estimate for suprema of Rademachers. It is a special case of [1, Lemma 5.10].
Proposition 10.
Let $T_{1},\ldots,T_{n}$ be nonempty bounded subsets of ${\mathbb{R}}^{N}$. Then
${\mathbb{E}}\max_{k\leq n}\sup_{t\in T_{k}}\sum_{i=1}^{N}t_{i}\varepsilon_{i}\lesssim\max_{k\leq n}{\mathbb{E}}\sup_{t\in T_{k}}\sum_{i=1}^{N}t_{i}\varepsilon_{i}+\max_{k\leq n}\sup_{t\in T_{k}}\Big{\|}\sum_{i=1}^{N}t_{i}\varepsilon_{i}\Big{\|}_{\mathrm{Log}\,n}.$
Another useful result is the estimate on the number of connected subsets of a graph.
Lemma 11.
Let $H=(V_{H},E_{H})$ be a graph with $n_{H}$ vertices and maximal degree $d_{H}$.
i) For a fixed $v\in V$ the number of connected subsets $I\subset V_{H}$ with cardinality $k$ containing $v$ is at most $(4d_{H})^{k-1}$.
ii) The number of all connected subsets $I\subset V_{H}$ with cardinality $k$ is not bigger than $n_{H}(4d_{H})^{k-1}$.
Proof.
i) The connected subset $I$ may be chosen by first choosing its spanning tree rooted at $v$ and then labelling the vertices of the tree. The number of unlabelled rooted trees is less than the number of oriented trees with $k$ vertices, i.e., less than the $(k-1)$-th Catalan number $C_{k-1}\leq 4^{k-1}$. The root of the tree is $v$ and the rest of vertices may be labelled in at most $d_{H}^{k-1}$ ways.
Part i) of the assertion immediately yields part ii). ∎
2.1. Van Handel-type bound
In this part we will establish the following improvement on van Handel’s bound [12].
Proposition 12.
For any $n\times m$ matrix $(a_{i,j})_{i\leq m,j\leq n}$ and $b\in(0,1]$ we have
${\mathbb{E}}\sup_{s\in B_{2}^{m}\cap bB_{\infty}^{m}}\sup_{t\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i\leq m,j\leq n}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$ $\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+\max_{j}\|(a_{i,j})_{i}\|_{2}$
$\phantom{\lesssim}+\mathrm{Log}((n+m)b^{2})\|(a_{i,j})_{i,j}\|_{\infty}.$
Let us first formulate and prove a symmetric variant of Propostion 12.
Proposition 13.
Let $(\tilde{\varepsilon}_{i,j})_{i,j}$ be a symmetric Rademacher matrix. Then for any symmetric matrix $(a_{i,j})_{i,j\leq n}$ and any $b\in(0,1]$,
${\mathbb{E}}\sup_{s,t\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i,j\leq n}a_{i,j}\tilde{\varepsilon}_{i,j}s_{i}t_{j}\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+\mathrm{Log}(nb^{2})\|(a_{i,j})_{i,j}\|_{\infty}.$
The proof uses the following, quite standard, technical lemma.
Lemma 14.
Let $Y_{1},\ldots,Y_{n}$ be r.v’s and $m_{i},\sigma_{i}\geq 0$ be such that
$\mathbb{P}(|Y_{i}|\geq m_{i}+u\sigma_{i})\leq e^{-u^{2}/2}\quad\mbox{ for every }u\geq 0\text{ and }i=1,\ldots,n.$
Then
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}^{2}Y_{i}\lesssim\max_{i}m_{i}+\sqrt{\mathrm{Log}(nb^{2})}\max_{i}\sigma_{i}$
and
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sqrt{\sum_{i=1}^{n}s_{i}^{2}Y_{i}^{2}}\lesssim\max_{i}m_{i}+\sqrt{\mathrm{Log}(nb^{2})}\max_{i}\sigma_{i}.$
Proof.
Let $(Y_{1}^{*},\ldots,Y_{n}^{*})$ be a nondecreasing rearrangement of $|Y_{1}|,\ldots,|Y_{n}|$. We set $k=n$ if $b^{2}\leq 1/n$, otherwise we choose $11\leq k\leq n-$ 1 such that $\frac{1}{k+1}<b^{2}\leq\frac{1}{k}$. Then $\mathrm{Log}(nb^{2})\sim\mathrm{Log}(n/k)$ and
$\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}^{2}Y_{i}\leq\frac{1}{k}(Y_{1}^{*}+\ldots+Y_{k}^{*}).$
A standard argument shows that ${\mathbb{E}}Y_{l}^{*}\leq({\mathbb{E}}|Y_{l}^{*}|^{2})^{1/2}\lesssim\max_{i}m_{i}+\mathrm{Log}^{1/2}(n/l)\max_{i}\sigma_{i}$. Thus
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}^{2}Y_{i}\lesssim\max_{i}m_{i}+\frac{1}{k}\sum_{l=1}^{k}\sqrt{\mathrm{Log}\bigl{(}\frac{n}{l}\bigr{)}}\max_{i}\sigma_{i}\lesssim\max_{i}m_{i}+\sqrt{\mathrm{Log}\bigl{(}\frac{n}{k}\bigr{)}}\max_{i}\sigma_{i}$
and
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sqrt{\sum_{i=1}^{n}s_{i}^{2}Y_{i}^{2}}$ $\leq{\mathbb{E}}\sqrt{\frac{1}{k}(|Y_{1}^{*}|^{2}+\ldots+|Y_{k}^{*}|^{2})}\leq\sqrt{\frac{1}{k}\sum_{l=1}^{k}{\mathbb{E}}|Y_{l}^{*}|^{2}}$
$\lesssim\max_{i}m_{i}+\sqrt{\frac{1}{k}\sum_{l=1}^{k}\mathrm{Log}\bigl{(}\frac{n}{l}\bigr{)}}\max_{i}\sigma_{i}$
$\lesssim\max_{i}m_{i}+\sqrt{\mathrm{Log}\bigl{(}\frac{n}{k}\bigr{)}}\max_{i}\sigma_{i}.\qed$
Proof of Proposition 13.
[…]
${\mathbb{E}}\|(a_{i,j}g_{i,j})\|={\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j}|g_{i,j}|)\|\geq{\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j}{\mathbb{E}}|g_{i,j}|)\|=\sqrt{\frac{2}{\pi}}{\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j})\|.$
For any symmetric matrix $B$ we have $\langle Bs,t\rangle=\frac{1}{4}\bigl{(}\langle B(s+t),s+t\rangle-\langle B(s-t),s-t\rangle\bigr{)}$, hence,
$\sup_{s,t\in B_{2}^{n}\cap bB_{\infty}^{n}}\langle Bs,t\rangle\leq 2\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}|\langle Bs,s\rangle|.$
Therefore,
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i,j\leq n}a_{i,j}\tilde{\varepsilon}_{i,j}s_{i}t_{j}$ $\lesssim{\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\Bigl{|}\sum_{i,j\leq n}a_{i,j}g_{i,j}s_{i}s_{j}\Bigr{|}$
$\leq{\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i,j\leq n}a_{i,j}g_{i,j}s_{i}s_{j}+{\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i,j\leq n}(-a_{i,j}g_{i,j}s_{i}s_{j})$
$=2{\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i,j\leq n}a_{i,j}g_{i,j}s_{i}s_{j}.$
Now we follow van Handel’s approach from [12]. Let $g_{1},g_{2},\ldots,g_{n}$ be iid $\mathcal{N}(0,1)$ r.v’s and $Y=(Y_{1},\ldots,Y_{n})\sim\mathcal{N}(0,B_{-})$, where $B_{-}$ is the negative part of $B=(a_{i,j}^{2})$. Define the new Gaussian process $Z_{s}$ by
$Z_{s}=2\sum_{i=1}^{n}s_{i}g_{i}\sqrt{\sum_{j=1}^{n}a_{ij}^{2}s_{j}^{2}}+\sum_{i=1}^{n}s_{i}^{2}Y_{i}.$
It is shown in [12] (see the proof of Theorem 4.1 therein) that for any $s,s^{\prime}\in{\mathbb{R}}^{n}$
${\mathbb{E}}\Bigl{|}\sum_{i,j\leq n}a_{i,j}g_{i,j}(s_{i}s_{j}-s_{i}^{\prime}s_{j}^{\prime})\Bigr{|}^{2}\leq{\mathbb{E}}|Z_{s}-Z_{s^{\prime}}|^{2}.$
Hence the Slepian-Fernique inequality [10, Theorem 3.15]. yields
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i,j\leq n}a_{i,j}g_{i,j}s_{i}s_{j}\leq{\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}Z_{s}.$
Variables $Y_{i}$ are centered Gaussian and (see the proof of Corollary 4.2 in [12]) $({\mathbb{E}}Y_{i}^{2})^{1/2}\leq\|(a_{i,j})_{j}\|_{4}$. Hence Lemma 14 applied with $m_{i}=0$ and $\sigma_{i}=\|(a_{i,j})_{j}\|_{4}$ yields
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}^{2}Y_{i}$ $\lesssim\sqrt{\mathrm{Log}(nb^{2})}\max_{i}\|(a_{i,j})_{j}\|_{4}$
$\leq\sqrt{\mathrm{Log}(nb^{2})}\max_{i,j}|a_{i,j}|^{1/2}\max_{j}\|(a_{i,j})_{j}\|_{2}^{1/2}$
$\leq\max_{i}\|(a_{i,j})_{j}\|_{2}+\mathrm{Log}(nb^{2})\|(a_{i,j})_{i,j}\|_{\infty}.$
We have
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}g_{i}\sqrt{\sum_{j}a_{ij}^{2}s_{j}^{2}}\leq{\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sqrt{\sum_{i,j}a_{ij}^{2}s_{j}^{2}g_{i}^{2}}={\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sqrt{\sum_{j}s_{j}^{2}V_{j}^{2}},$
where $V_{j}=\sqrt{\sum_{i}a_{ij}^{2}g_{i}^{2}}$. The Gaussian concentration [10, Lemma 3.1] yields
$\mathbb{P}(|V_{j}|\geq\|(a_{i,j})_{i}\|_{2}+t\|(a_{i,j})_{i}\|_{\infty})\leq e^{-t^{2}/2},$
so Lemma 14 applied with $Y_{j}=V_{j}$, $m_{j}=\|(a_{i,j})_{i}\|_{2}$ and $\sigma_{j}=\|(a_{i,j})_{i}\|_{\infty}$ yields
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}\sqrt{\sum_{j}a_{ij}^{2}s_{j}^{2}}g_{i}\lesssim\max_{j}\|(a_{i,j})_{i}\|_{2}+\sqrt{\mathrm{Log}(nb^{2})}\|(a_{i,j})_{i,j}\|_{\infty}.\qed$
Proof of Proposition 12.
We apply Proposition 13 to the symmetric $(n+m)\times(n+m)$ matrix $\tilde{A}=(\tilde{a}_{i,j})$ of the form $\tilde{A}=\left(\begin{array}[]{cc}0&A\\ A^{T}&0\end{array}\right)$. Observe that $\max_{i,j}|\tilde{a}_{i,j}|=\max_{i,j}|a_{i,j}|$,
$\max_{i}\|(\tilde{a}_{i,j})_{j}\|_{2}=\max\Bigl{\{}\max_{i}\|(a_{i,j})_{j}\|_{2},\max_{j}\|(a_{i,j})_{i}\|_{2}\Bigr{\}}$
and
${\mathbb{E}}\sup_{s,t\in B_{2}^{n+m}\cap bB_{\infty}^{n+m}}$ $\sum_{i,j\leq n+m}\tilde{a}_{i,j}\tilde{\varepsilon}_{i,j}s_{i}t_{j}\geq{\mathbb{E}}\sup_{s\in B_{2}^{m}\cap bB_{\infty}^{m}}\sup_{t\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i\leq m,j\leq n}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.\qed$
3. Bounds up to polylog factors
Let us first introduce the notation which will be used till the end of the paper. Recall that $G_{A}=([n],E_{A})$ is a graph associated to a given symmetric matrix $A=(a_{i,j})_{i,j\leq n}$. By $\rho=\rho_{A}$ we denote the distance on $[n]$ induced by $E_{A}$. For $r=1,2,\ldots$ we put $G_{r}=G_{r}(A)=([n],E_{A,r})$, where $(i,j)\in E_{A,r}$ iff $\rho(i,j)\leq r$. In particular $G_{1}=([n],E_{A})$ and the maximal degree of $G_{r}$ is at most $d_{A}+d_{A}(d_{A}-1)+\ldots+d_{A}(d_{A}-1)^{r-1}\leq d_{A}^{r}$. We say that a subset of $[n]$ is $r$-connected if it is connected in $G_{r}$.
We denote by $\mathcal{I}(k)=\mathcal{I}(k,n)$ the family of all subsets of $[n]$ of cardinality $k$ and by $\mathcal{I}_{r}(k)=\mathcal{I}_{r}(k,A)$ the family of all $r$-connected subsets of $[n]$ of cardinality $k$.
For a set $I\subset[n]$ and a vertex $j\in[n]$ we write $I\sim_{A}j$ if $(i,j)\in E_{A}$ for some $i\in I$. By $I^{\prime}=I^{\prime}(A)$ we denote the set of all neighbours of $I$ in $G_{1}$ and by $I^{\prime\prime}=I^{\prime\prime}(A)$ the set of all neighbours of $I^{\prime}$ in $G_{1}$, i.e.,
$I^{\prime}=\{j\in[n]\colon\ \exists_{i\in I}\ (i,j)\in E_{A}\},\quad I^{\prime\prime}=\{i\in[n]\colon\ \exists_{i_{0}\in I,j\in[n]}\ (i_{0},j),(i,j)\in E_{A}\}.$ (10)
Observe that $I$ is a subset of $I^{\prime\prime}$, but does not have to be a subset of $I^{\prime}$. Moreover $|I^{\prime}|\leq d_{A}|I|$ and $|I^{\prime\prime}|\leq d_{A}^{2}|I|$.
By Remark 6 we may and will assume that $a_{i,i}=0$ for all $i$.
For $1\leq k,l\leq n$ define random variables
$X_{k,l}=X_{k,l}(A):=\frac{1}{\sqrt{kl}}\max_{I\in\mathcal{I}(k),J\in\mathcal{I}(l)}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}^{\prime}$
and their $4$-connected counteparts
$\overline{X}_{k,l}=\overline{X}_{k,l}(A):=\frac{1}{\sqrt{kl}}\max_{I\in\mathcal{I}_{4}(k),J\in\mathcal{I}_{4}(l)}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}^{\prime}.$
Set
$X=X(A):=\max_{1\leq k,l\leq n}X_{k,l}=\max_{\emptyset\neq I,J\subset V}\frac{1}{\sqrt{|I||J|}}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}.$ (11)
Lemma 15.
For any $1\leq k,l\leq n$,
${\mathbb{E}}X_{k,l}\lesssim\max_{1\leq k^{\prime}\leq k,1\leq l^{\prime}\leq l}{\mathbb{E}}\overline{X}_{k^{\prime},l^{\prime}}+R_{A}(\mathrm{Log}(kl)).$
Proof.
Let us first fix sets $I\in\mathcal{I}(k)$ and $J\in\mathcal{I}(l)$. Let $I_{1},\ldots,I_{r}$ be connected components of $I\cap J^{\prime}$ in $G_{2}$ and $J_{u}:=J\cap I_{u}^{\prime}$. Then sets $J_{1},\ldots,J_{r}$ are disjoint and 4-connected subsets of $J$. Hence, for every $\eta_{i},\eta_{j}^{\prime}=\pm 1$ we have
$\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}^{\prime}$ $=\sum_{i\in I\cap J^{\prime},j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}^{\prime}=\sum_{u=1}^{r}\sum_{i\in I_{u},j\in J_{u}}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}^{\prime}\leq\sum_{u=1}^{r}\overline{X}_{|I_{u}|,|J_{u}|}\sqrt{|I_{u}||J_{u}|}$
$\leq\max_{k^{\prime}\leq k,l^{\prime}\leq l}\overline{X}_{k^{\prime},l^{\prime}}\sum_{u=1}^{r}\sqrt{|I_{u}||J_{u}|}\leq\max_{k^{\prime}\leq k,l^{\prime}\leq l}\overline{X}_{k^{\prime},l^{\prime}}\Bigl{(}\sum_{u=1}^{r}|I_{u}|\Bigr{)}^{1/2}\Bigl{(}\sum_{u=1}^{r}|J_{u}|\Bigr{)}^{1/2}$
$\leq\max_{k^{\prime}\leq k,l^{\prime}\leq l}\overline{X}_{k^{\prime},l^{\prime}}\sqrt{|I||J|}.$
Taking the supremum over all sets $I\in\mathcal{I}(k)$, $J\in\mathcal{I}(l)$ and $\eta_{i},\eta_{j}^{\prime}=\pm 1$ we get
$X_{k,l}\leq\max_{k^{\prime}\leq k,l^{\prime}\leq l}\overline{X}_{k^{\prime},l^{\prime}}.$ (12)
Observe that
$\max_{I\in\mathcal{I}_{4}(k^{\prime}),J\in\mathcal{I}_{4}(l^{\prime})}\max_{\eta_{i},\eta^{\prime}_{j}=\pm 1}\frac{1}{\sqrt{k^{\prime}l^{\prime}}}\Bigl{\|}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}\Bigr{\|}_{\mathrm{Log}(kl)}$ $\leq\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\Bigl{\|}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{\|}_{\mathrm{Log}(kl)}$
$=R_{A}(\mathrm{Log}(kl)).$
Thus, by Proposition 10
${\mathbb{E}}\max_{k^{\prime}\leq k,l^{\prime}\leq l}\overline{X}_{k^{\prime},l^{\prime}}\lesssim\max_{k^{\prime}\leq k,l^{\prime}\leq l}{\mathbb{E}}\overline{X}_{k^{\prime},l^{\prime}}+R_{A}(\mathrm{Log}(kl)).$
Lemma 16.
We have for any $1\leq k,l\leq n$,
${\mathbb{E}}\overline{X}_{k,l}\lesssim\sqrt{\mathrm{Log}\,d_{A}}\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n).$
Proof.
Obviously $\overline{X}_{k^{\prime},l^{\prime}}\leq\|(a_{i,j}\varepsilon_{i,j})\|$, so by (7) we may assume that $n\geq d_{A}\geq 3$. By the symmetry it is enough to consider only the case $l\geq k$. By Lemma 11, $2^{k}|\mathcal{I}_{4}(k)|\leq n(8d_{A}^{4})^{k}\leq nd_{A}^{6k}$.
We have
$\overline{X}_{k,l}\leq\frac{1}{\sqrt{k}}\max_{I\in\mathcal{I}_{4}(k)}\max_{\eta_{i}=\pm 1}\sup_{\|t\|_{2}\leq 1}\sum_{i\in I}\sum_{j}a_{i,j}\varepsilon_{i,j}\eta_{i}t_{j}.$
For any fixed $I\in\mathcal{I}_{4}(k)$ and $\eta_{i}=\pm 1$,
${\mathbb{E}}\frac{1}{\sqrt{k}}\sup_{\|t\|_{2}\leq 1}\sum_{i\in I}\sum_{j}a_{i,j}\varepsilon_{i,j}\eta_{i}t_{j}$ $=\frac{1}{\sqrt{k}}{\mathbb{E}}\Bigl{(}\sum_{j}\Bigr{(}\sum_{i\in I}a_{i,j}\varepsilon_{i,j}\eta_{i}\Bigl{)}^{2}\Bigr{)}^{1/2}$
$\leq\frac{1}{\sqrt{k}}\Bigl{(}\sum_{j}{\mathbb{E}}\Bigr{(}\sum_{i\in I}a_{i,j}\varepsilon_{i,j}\eta_{i}\Bigl{)}^{2}\Bigr{)}^{1/2}=\frac{1}{\sqrt{k}}\Bigl{(}\sum_{j}\sum_{i\in I}a_{i,j}^{2}\Bigr{)}^{1/2}$
$=\frac{1}{\sqrt{k}}\Bigl{(}\sum_{i\in I}\sum_{j}a_{i,j}^{2}\Bigr{)}^{1/2}\leq\max_{i}\|(a_{i,j})_{j}\|_{2}.$
In the case $n\geq d_{A}^{6k}$, $\log(2^{k}|\mathcal{I}_{4}(k)|)\lesssim\log n$ and
$\max_{I\in\mathcal{I}_{4}(k)}\max_{\eta_{i}=\pm 1}\sup_{\|t\|_{2}\leq 1}\frac{1}{\sqrt{k}}\Bigr{\|}\sum_{i\in I}\sum_{j}a_{i,j}\varepsilon_{i,j}\eta_{i}t_{j}\Bigl{\|}_{\mathrm{Log}(2^{k}|\mathcal{I}_{4}(k)|)}\lesssim R_{A}(\log n).$
In the case $n\leq d_{A}^{6k}$ we have $\log(2^{k}|\mathcal{I}_{4}(k)|)\lesssim k\log d_{A}$ and
$\max_{I\in\mathcal{I}_{4}(k)}$ $\max_{\eta_{i}=\pm 1}\sup_{\|t\|_{2}\leq 1}\frac{1}{\sqrt{k}}\Bigr{\|}\sum_{i\in I}\sum_{j}a_{i,j}\varepsilon_{i,j}\eta_{i}t_{j}\Bigl{\|}_{\mathrm{Log}(2^{k}|\mathcal{I}_{4}(k)|)}$
$\lesssim\max_{I\in\mathcal{I}_{4}(k)}\max_{\eta_{i}=\pm 1}\sup_{\|t\|_{2}\leq 1}\sqrt{\log d_{A}}\Bigl{(}\sum_{i\in I}\sum_{j}(a_{i,j}\eta_{i}t_{j})^{2}\Bigr{)}^{1/2}$
$=\max_{I\in\mathcal{I}_{4}(k)}\max_{j}\sqrt{\log d_{A}}\Bigl{(}\sum_{i\in I}a_{i,j}^{2}\Bigr{)}^{1/2}\leq\sqrt{\log d_{A}}\max_{j}\|(a_{i,j})_{i}\|_{2}.$
The assertion follows by Proposition 10. ∎
Corollary 17.
We have
${\mathbb{E}}X={\mathbb{E}}\max_{\emptyset\neq I,J\subset V}\frac{1}{\sqrt{|I||J|}}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}\lesssim\sqrt{\mathrm{Log}\,d_{A}}\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n).$
Proof.
Lemmas 15 and 16 imply that for a fixed $1\leq k,l\leq n$
${\mathbb{E}}X_{k,l}\lesssim\max_{k^{\prime}\leq k,l^{\prime}\leq l}{\mathbb{E}}\overline{X}_{k^{\prime},l^{\prime}}+R_{A}(\mathrm{Log}(kl))\lesssim\sqrt{\mathrm{Log}\,d_{A}}\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n).$
Moreover,
$\max_{1\leq k,l\leq n}\max_{I\in\mathcal{I}(k),J\in\mathcal{I}(l)}\frac{1}{\sqrt{kl}}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}$ $\Big{\|}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}\Big{\|}_{\mathrm{Log}(n^{2})}$
$\leq\sup_{\|t\|_{2}\leq 1,\|s\|_{2}\leq 1}\Big{\|}\sum_{i,j\in V}a_{i,j}\varepsilon_{i,j}t_{i}s_{j}\Big{\|}_{2\mathrm{Log}\,n}\lesssim R_{A}(\mathrm{Log}\,n)$
and the assertion follows by Proposition 10. ∎
Proposition 18.
For any symmeric matrix $(a_{i,j})_{i,j\leq n}$ we have
${\mathbb{E}}\bigl{\|}(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\bigr{\|}\lesssim\mathrm{Log}^{3/2}(d_{A})\max_{i}\|(a_{i,j})_{j}\|_{2}+\mathrm{Log}(d_{A})R_{A}(\mathrm{Log}\,n).$
Proof.
By Remark 6 we may assume that $a_{i,i}=0$ for all $i$ and $n\geq d_{A}\geq 3$.
For vectors $s,t$ and integers $k,l$ we define sets
$I_{k}(s)=\{i\in V\colon\ e^{-k-1}<|s_{i}|\leq e^{-k}\},\quad J_{l}(t)=\{j\in V\colon\ e^{-l-1}<|t_{j}|\leq e^{-l}\}.$
Observe that for any $s,t,k,l$,
$\sum_{i\in I_{k}(s),j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\leq e^{-k-l}\max_{\eta_{i},\eta^{\prime}_{j}=\pm 1}\sum_{i\in I_{k}(s),j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j},$
therefore
$\bigl{\|}(a_{i,j}\varepsilon_{i,j})_{i,j\in V}\bigr{\|}\leq\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\sum_{k,l}e^{-k-l}\sup_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I_{k}(s),j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}.$
We have
$\sum_{k}\sum_{l\geq k+\log d_{A}}$ $e^{-k-l}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I_{k}(s),j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}$
$\leq\sum_{k}\sum_{l\geq k+\log d_{A}}e^{-k}\sum_{i\in I_{k}(s),j\in J_{l}(t)}|a_{i,j}|e^{-l}$
$\leq\sum_{k}e^{-k}\sum_{i\in I_{k}(s)}\sum_{j}|a_{i,j}|\sum_{l\geq k+\log d_{A}}e^{-l}\mathbbm{1}_{\{j\in J_{l}(t)\}}$
$\leq\sum_{k}\sum_{i\in I_{k}(s)}\sum_{j}|a_{i,j}|e^{-2k-\log d_{A}}\leq\|(a_{i,j})\|_{\infty}\sum_{k}\sum_{i\in I_{k}(s)}e^{-2k}$
$\leq\|(a_{i,j})\|_{\infty}\sum_{k}\sum_{i\in I_{k}(s)}e^{2}s_{i}^{2}=e^{2}\|s\|_{2}^{2}\|(a_{i,j})\|_{\infty}.$
In the same way we show that
$\sum_{l}\sum_{k\geq l+\log d_{A}}e^{-k-l}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I_{k}(s),j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}\leq e^{2}\|t\|_{2}^{2}\|(a_{i,j})\|_{\infty}.$
Moreover, for any $s,t$,
$\sum_{k,l\colon|k-l|<\log d_{A}}e^{-k-l}\sup_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I_{k}(s),j\in J_{l}(t)}$ $a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}$
$\leq X\sum_{k,l\colon|k-l|<\log d_{A}}^{\infty}e^{-k-l}\sqrt{|I_{k}(s)||J_{l}(t)|}.$
For any fixed integer $r$
$\sum_{k}e^{-k-(k+r)}\sqrt{|I_{k}(s)||J_{k+r}(t)|}$ $\leq\Bigl{(}\sum_{k}e^{-2k}|I_{k}(s)|\Bigr{)}^{1/2}\Bigl{(}\sum_{k}e^{-2(k+r)}|J_{k+r}(t)|\Bigr{)}^{1/2}$
$\leq e^{2}\|t\|_{2}\|s\|_{2}.$
Hence,
$\bigl{\|}(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\bigr{\|}$ $\leq\sup_{\|s\|_{2},\|t\|_{2}\leq 1}e^{2}((\|s\|^{2}+\|t\|^{2})\|(a_{i,j})\|_{\infty}+(2\log d_{A}+1)X\|t\|_{2}\|s\|_{2})$
$\leq e^{2}(2\|(a_{i,j})\|_{\infty}+(2\log d_{A}+1)X)$
and the assertion follows by Corollary 17. ∎
4. Proof of Theorem 5
By Remark 6 we may assume that $a_{i,i}=0$ for all $i$ and $n\geq d_{A}\geq 3$.
For $k=1,2,\ldots$ and $t,s\in B_{2}^{n}$ we define
$I_{k}(s):=\{i\colon\ d_{A}^{-k/40}<|s_{i}|\leq d_{A}^{(1-k)/40}\},\quad J_{l}(t):=\{j\colon\ d_{A}^{-l/40}<|t_{j}|\leq d_{A}^{(1-l)/40}\}.$
Then
$\sum_{k\geq 1}d_{A}^{-k/20}|I_{k}(s)|\leq\|s\|_{2}^{2},\quad\sum_{l\geq 1}d_{A}^{-l/20}|J_{l}(t)|\leq\|t\|_{2}^{2}$ (13)
and
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|={\mathbb{E}}\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k,l\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$
Observe that for any $s,t\in B_{2}^{n}$,
$\Bigl{|}\sum_{k\geq 1}\sum_{l\geq k+41}\sum_{i\in I_{k}(s)}\sum_{j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{|}$ $\leq\sum_{k\geq 1}\sum_{i\in I_{k}(s)}|s_{i}|\sum_{l\geq k+41}\sum_{j\in J_{l}(t)}|a_{i,j}||t_{j}|$
$\leq\sum_{k\geq 1}\sum_{i\in I_{k}(s)}|s_{i}|d_{A}^{-(k+40)/40}\sum_{j}|a_{i,j}|$
$\leq\|(a_{i,j})\|_{\infty}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}s_{i}^{2}\leq\|(a_{i,j})\|_{\infty}.$
Similarily,
$\Bigl{|}\sum_{l\geq 1}\sum_{k\geq l+41}\sum_{i\in I_{k}(s)}\sum_{j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{|}\leq\|(a_{i,j})\|_{\infty}.$
Hence it is enough to estimate
$\sum_{r=-40}^{40}{\mathbb{E}}\sup_{\|s\|_{2}\leq}\sup_{\|t\|_{2}\leq 1}\sum_{\begin{subarray}{c}k,l\geq 1\\ l-k=r\end{subarray}}\sum_{i\in I_{k}(s)}\sum_{j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$
By symmetry it is enough to bound only the terms with $r\geq 0$. Let $X$ be defined by (11).
Then for a fixed $r\geq 0$ and $\alpha>0$,
$\sup_{\|s\|_{2}\leq 1}$ $\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\{|J_{k+r}(t)|\geq\alpha|I_{k}(s)|\}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\leq\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}d_{A}^{(2-2k-r)/40}\mathbbm{1}_{\{|J_{k+r}(t)|\geq\alpha|I_{k}(s)|\}}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}^{\prime}$
$\leq X\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}d_{A}^{(2-2k-r)/40}\sqrt{|I_{k}(s)||J_{k+r}(t)|}\mathbbm{1}_{\{|J_{k+r}(t)|\geq\alpha|I_{k}(s)|\}}$
$\leq\alpha^{-1/2}X\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}d_{A}^{(2-2k-r)/40}|J_{k+r}(t)|\leq\alpha^{-1/2}d_{A}^{(r+2)/40}X,$
where the last inequality follows by (13).
Hence Corollary 17 yields
${\mathbb{E}}\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\{|J_{k+r}(t)|\geq d_{A}^{(2r+5)/40}|I_{k}(s)|\}}$ $a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\log n).$
In a similar way we show that
$\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}$ $\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\{|J_{k+r}(t)|\leq\alpha|I_{k}(s)|\}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\leq\alpha^{1/2}X\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}d_{A}^{(2-2k-r)/40}|I_{k}(s)|\leq\alpha^{1/2}d_{A}^{(2-r)/40}X$
and
${\mathbb{E}}\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\{|J_{k+r}(t)|\geq d_{A}^{(2r-5)/40}|I_{k}(s)|\}}$ $a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\log n).$
Hence it is enough to bound for $r=0,1,\ldots,40$,
${\mathbb{E}}\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\left\{d_{A}^{(2r-5)/40}<|J_{k+r}(t)|/|I_{k}(s)|<d_{A}^{(2r+5)/40}\right\}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$
Recall definition (10) of sets $I^{\prime}$ and $I^{\prime\prime}$. Let us fix $0\leq r\leq 40$ and $k,t,s$ such that $d_{A}^{(2r-5)/40}<|J_{k+r}(t)|/|I_{k}(s)|<d_{A}^{(2r+5)/40}$. Let $|I_{k}(s)|=m$. Let us consider the following greedy algorithm with output being a subset $\{i_{1},\ldots,i_{M}\}$ of $I_{k}(s)$ of size $M\leq m$
  • In the first step we pick a vertex $i_{1}\in I_{k}(s)$ with maximal number of neighbours in $J_{k+r}(t)$.
  • Once we have $\{i_{1},\ldots,i_{N}\}$ and $N<M$, we pick $i_{N+1}\in I_{k}(s)\setminus\{i_{1},\ldots,i_{N}\}$ with maximal number of neighbours in $J_{k+r}(t)\setminus\{i_{1},\ldots,i_{N}\}^{\prime}$.
If $l_{N}$ is the number of neighbours of $i_{N}$ in $J_{k+r}(t)\setminus\{i_{1},\ldots,i_{N-1}\}^{\prime}$, then $l_{1}\geq l_{2}\geq\ldots\geq l_{M}$, so $Ml_{M}\leq|J_{k+r}(t)|$. Hence, using this algorithm we may find a subset $I\subset I_{k}(s)$ with cardinality $|I|\leq d_{A}^{-(r+18)/40}|J_{k+r}(t)|\leq d_{A}^{(r-13)/40}m$ such that for every $i\in I_{k}(s)\setminus I$, |$|\{j\in J_{k+r}(t)\setminus I^{\prime}\colon i\sim_{A}j\}|\leq\lceil d_{A}^{(r+18)/40}\rceil\leq 2d_{A}^{(r+18)/40}$. Note that if $(i,j)\in E_{A}$ and $(i,j)\in(I_{k}(s)\times J_{k+r}(t))\setminus(I^{\prime\prime}\times I^{\prime})$, then $j\in J_{k+r}(t)\setminus I^{\prime}$. Therefore,
$\sum_{(i,j)\in I_{k}(s)\times J_{k+r}(t))\setminus(I^{\prime\prime}\times I^{\prime})}|a_{i,j}||s_{i}t_{j}|$ $\leq 2\|(a_{i,j})\|_{\infty}\sum_{i\in I_{k}(s)}|s_{i}|d_{A}^{(r+18)/40}d_{A}^{(1-k-r)/40}$
$\leq 2\|(a_{i,j})\|_{\infty}d_{A}^{19/40}\sum_{i\in I_{k}(s)}s_{i}^{2}.$ (14)
Observe that if $d_{A}^{-(r+18)/40}|J_{k+r}(t)|>m$ we may take $I=I_{k}(s)$ and then $a_{i,j}=0$ for $(i,j)\in I_{k}(s)\times J_{k+r}(t))\setminus(I^{\prime\prime}\times I^{\prime})$, so estimate (14) is also valid in this case.
Let
$s^{\prime}=(s^{\prime}_{i})_{i\in I^{\prime\prime}\cap I_{k}(s)},\quad t^{\prime}=(t^{\prime}_{j})_{j\in I^{\prime}\cap J_{k+r}(s)},$
where
$s^{\prime}_{i}:=\frac{s_{i}}{\|(s_{i})_{i\in I_{k}(s)}\|_{2}},\quad t^{\prime}_{j}:=\frac{t_{j}}{\|(t_{j})_{j\in J_{k+r}(t)}\|_{2}}.$
Then
$\|s^{\prime}\|_{2}\leq 1,\quad\|s^{\prime}\|_{\infty}\leq d_{A}^{1/40}|I_{k}(s)|^{-1/2}=d_{A}^{1/40}m^{-1/2},$
$\|t^{\prime}\|_{2}\leq 1,\quad\|t^{\prime}\|_{\infty}\leq d_{A}^{1/40}|J_{k+r}(t)|^{-1/2}\leq d_{A}^{(7-2r)/80}m^{-1/2}.$
Hence,
$\sum_{(i,j)\in(I_{k}(s)\times J_{k+r}(t))\cap(I^{\prime\prime}\times I^{\prime})}\varepsilon_{i,j}a_{i,j}s_{i}t_{j}\leq Y_{m,r}\|(s_{i})_{i\in I_{k}(s)}\|_{2}\|(t_{j})_{j\in J_{k+r}(t)}\|_{2},$ (15)
where
$Y_{m,r}:=\max_{|I|\leq d_{A}^{(r-13)/40}m}\sup_{s\in B_{2}^{n}\cap d_{A}^{1/40}m^{-1/2}B_{\infty}^{n}}\ \sup_{t\in B_{2}^{n}\cap d_{A}^{(7-2r)/80}m^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$
Define $Y_{r}:=\max_{1\leq m\leq n}Y_{m,r}$. Estimates (14) and (15) yield
${\mathbb{E}}$ $\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\Big{\{}d_{A}^{(2r-5)/40}<|J_{k+r}(t)|/|I_{k}(s)|<d_{A}^{(2r+5)/40}\Big{\}}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\leq\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\Biggl{(}\sum_{k\geq 1}2d_{A}^{19/40}\|(a_{i,j})\|_{\infty}\sum_{i\in I_{k}(s)}s_{i}^{2}+{\mathbb{E}}Y_{r}\sum_{k\geq 1}\|(s_{i})_{i\in I_{k}(s)}\|_{2}\|(t_{j})_{j\in J_{k+r}(t)}\|_{2}\Biggr{)}$
$\leq 2d_{A}^{19/40}\|(a_{i,j})\|_{\infty}+{\mathbb{E}}Y_{r}\sup_{\|s\|_{2}\leq 1}\Bigl{(}\sum_{k\geq 1}\|(s_{i})_{i\in I_{k}(s)}\|_{2}^{2}\Bigr{)}^{1/2}\sup_{\|t\|_{2}\leq 1}\Bigl{(}\sum_{k\geq 1}\|(t_{j})_{j\in J_{k+r}(t)}\|_{2}^{2}\Bigr{)}^{1/2}$
$\leq 2d_{A}^{19/40}\|(a_{i,j})\|_{\infty}+{\mathbb{E}}Y_{r}.$
Therefore, to establish Theorem 5 it is enough to prove the following lemma.
Lemma 19.
For every $0\leq r\leq 40$,
${\mathbb{E}}Y_{r}={\mathbb{E}}\max_{1\leq m\leq n}Y_{m,r}\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\log n)+\mathrm{Log}(d_{A})\|(a_{i,j})\|_{\infty}.$
First we show a connected counterpart to Lemma 19.
Lemma 20.
We have
${\mathbb{E}}\max_{1\leq k\leq n}\max_{I\in\mathcal{I}_{4}(k)}\ \sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}$ $\sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\log n)+\mathrm{Log}(d_{A})\|(a_{i,j})\|_{\infty}.$
Proof.
Let us first fix $k$ and $I\in\mathcal{I}_{4}(k)$. Then $|I^{\prime}|\leq d_{A}k$ and $|I^{\prime\prime}|\leq d_{A}^{2}k$. Proposition 12, applied with $(a_{i,j})=(a_{i,j})_{i\in I^{\prime\prime},j\in I^{\prime}}$, $n=|I^{\prime\prime}|$, $m=|I^{\prime}|$, and $b=d_{A}^{3/8}k^{-1/2}$ yields
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}\ \sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}$ $a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+\mathrm{Log}(d_{A})\|(a_{i,j})\|_{\infty}.$
By Lemma 11, $|\mathcal{I}_{4}(k)|\leq n(4d_{A}^{4})^{k}\leq\max\{n^{2},d_{A}^{12k}\}$ (recall that we assume that $d_{A}\geq 3$). We have
$\sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}\ \sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\Bigl{\|}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{\|}_{\mathrm{Log}(|\mathcal{I}_{4}(k)|)}$
$\leq\sup_{s,t\in B_{2}^{n}}\Bigl{\|}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{\|}_{2\mathrm{Log}(n)}+\sup_{s\in B_{2}^{n}}\sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\Bigl{\|}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{\|}_{12k\mathrm{Log}(d_{A})}$
$\leq R_{A}(2\mathrm{Log}n)+\sup_{s\in B_{2}^{n}}\sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\sqrt{12k\mathrm{Log}(d_{A})}\Bigl{(}\sum_{i,j}a_{i,j}^{2}s_{i}^{2}t_{j}^{2}\Bigr{)}^{1/2}$
$\lesssim R_{A}(\mathrm{Log}\,n)+\sqrt{k\mathrm{Log}(d_{A})}\max_{i}d_{A}^{-1/16}k^{-1/2}\Bigl{(}\sum_{j}a_{i,j}^{2}\Bigr{)}^{1/2}$
$\leq R_{A}(\mathrm{Log}\,n)+\sqrt{\mathrm{Log}(d_{A})}d_{A}^{-1/16}\max_{i}\|(a_{i,j})_{j}\|_{2}\lesssim R_{A}(\log n)+\max_{i}\|(a_{i,j})_{j}\|_{2}.$
Hence, by Proposition 10,
${\mathbb{E}}\max_{I\in\mathcal{I}_{4}(k)}\sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}$ $\sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\log n)+\mathrm{Log}(d_{A})\|(a_{i,j})\|_{\infty}.$
Applying again Proposition 10 and observing that
$\max_{1\leq k\leq n}\sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}\ \sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \Bigl{\|}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{\|}_{\mathrm{Log}(n)}\leq R_{A}(\log n)$
we get the assertion. ∎
Proof of Lemma 19.
Let
$Z_{k}:=\max_{I\in\mathcal{I}_{4}(k)}\sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}\ \sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$
Let us fix $r\geq 0$, $I\subset V$ such that $|I|\leq d_{A}^{(r-13)/40}m$, $s\in B_{2}^{n}\cap d_{A}^{1/40}m^{-1/2}B_{\infty}^{n}$ and $t\in B_{2}^{n}\cap d_{A}^{(7-2r)/80}m^{-1/2}B_{\infty}^{n}$. Let $I_{1},\ldots,I_{l}$ be 4-connected components of $I$. Then $\{I_{1}^{\prime},\ldots,I_{l}^{\prime}\}$ is a partition of $I^{\prime}$, $\{I_{1}^{\prime\prime},\ldots,I_{l}^{\prime\prime}\}$ is a partition of $I^{\prime\prime}$ and
$\sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}=\sum_{u=1}^{l}\sum_{i\in I_{u}^{\prime\prime},j\in I_{u}^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$ (16)
Define
$U_{1}$ :=$:=\bigl{\{}1\leq u\leq l\colon\ \|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}\geq d_{A}^{(12-r)/80}m^{-1/2}\sqrt{|I_{u}|}\bigr{\}},$
$U_{2}$ :=$:=\bigl{\{}1\leq u\leq l\colon\ \|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}\geq d_{A}^{(12-r)/80}m^{-1/2}\sqrt{|I_{u}|}\bigr{\}}.$
For $u\in U_{1}\cap U_{2}$ define vectors
$\tilde{s}(u):=\frac{(s_{i})_{i\in I_{u}^{\prime\prime}}}{\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}},\quad\tilde{t}(u):=\frac{(t_{j})_{j\in I_{u}^{\prime}}}{\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}}.$
Then $\|\tilde{s}(u)\|_{2}=\|\tilde{t}(u)\|_{2}=1$,
$\|\tilde{s}(u)\|_{\infty}$ $\leq d_{A}^{(r-12)/80}m^{1/2}|I_{u}|^{-1/2}\|s\|_{\infty}\leq d_{A}^{(r-10)/80}|I_{u}|^{-1/2}\leq d_{A}^{3/8}|I_{u}|^{-1/2},$
$\|\tilde{t}(u)\|_{\infty}$ $\leq d_{A}^{(r-12)/80}m^{1/2}|I_{u}|^{-1/2}\|t\|_{\infty}\leq d_{A}^{-(r+5)/80}|I_{u}|^{-1/2}\leq d_{A}^{-1/16}|I_{u}|^{-1/2}.$
Hence
$\sum_{u\in U_{1}\cap U_{2}}\sum_{i\in I_{u}^{\prime\prime},j\in I_{u}^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$ $\leq\sum_{u\in U_{1}\cap U_{2}}Z_{|I_{u}|}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}$
$\leq\max_{k}Z_{k}\Bigl{(}\sum_{u\leq l}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}^{2}\Bigr{)}^{1/2}\Bigl{(}\sum_{u\leq l}\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}^{2}\Bigr{)}^{1/2}$
$\leq\max_{k}Z_{k}.$ (17)
Observe that
$\sum_{u\notin U_{1}}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}^{2}\leq\sum_{u}d_{A}^{(12-r)/40}m^{-1}|I_{u}|=d_{A}^{(12-r)/40}m^{-1}|I|\leq d_{A}^{-1/40}$
and by the same token
$\sum_{u\notin U_{2}}\|(t_{j})_{i\in I_{u}^{\prime}}\|_{2}^{2}\leq d_{A}^{-1/40}.$
Hence
$\sum_{u\notin U_{1}}\sum_{i\in I_{u}^{\prime\prime},j\in I_{u}^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$ $\leq\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|\sum_{u\notin U_{1}}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}$
$\leq\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|\Bigl{(}\sum_{u\notin U_{1}}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}^{2}\Bigr{)}^{1/2}\Bigl{(}\sum_{u\leq l}\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}^{2}\Bigr{)}^{1/2}$
$\leq d_{A}^{-1/80}\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|$ (18)
and
$\sum_{u\in U_{1}\setminus U_{2}}\sum_{i\in I_{u}^{\prime\prime},j\in I_{u}^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$ $\leq\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|\sum_{u\notin U_{2}}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}$
$\leq\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|\Bigl{(}\sum_{u\leq l}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}^{2}\Bigr{)}^{1/2}\Bigl{(}\sum_{u\notin U_{2}}\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}^{2}\Bigr{)}^{1/2}$
$\leq d_{A}^{-1/80}\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|.$ (19)
Bounds (16)-(19) yield
${\mathbb{E}}\max_{m}Y_{m,r}\leq{\mathbb{E}}\max_{k}Z_{k}+2d_{A}^{-1/80}{\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|$
and the assertion follows by Lemma 20 and Proposition 18. ∎
Acknowledgements. Part of this work was carried out while the author was visiting the Hausdorff Research Institute for Mathematics, Univerity of Bonn. The hospitality of HIM and of the organizers of the program Synergies between modern probability, geometric analysis and stochastic geometry is gratefully acknowledged. The author would like to thank Marta Strzelecka for careful readings of various versions of the manuscript and numerous valuable comments and Ramon van Handel for pointing out the reference [6].