### v2

# On the spectral norm of Rademacher matrices

Abstract.
We discuss two-sided non-asymptotic bounds for the mean spectral norm of nonhomogenous weighted Rademacher matrices. We show that the recently formulated conjecture holds up to $\log\log\log n$ factor for arbitrary $n\times n$ Rademacher matrices and the triple logarithm may be eliminated for matrices with $\{0,1\}$-coefficients.
2020 Mathematics Subject Classification:
Primary 60B20, Secondary 15B52, 46B09
1. Introduction and main results
One of the basic issues of the random matrix theory are bounds on the spectral norm (largest singular value) of various families of random matrices. This question is very well understood for classical ensembles of random matrices [3], when one may use methods based on the large degree of symmetry. Recently, a substantial progress was attained in the understanding of inhomogeneous models [18], especially in the Gaussian case [12, 4]. However, there are still many open questions in this area, the one concerning Rademacher matrices is discussed here.
In this paper we investigate the mean operator (spectral) norm of weighted Rademacher matrices, i.e., quantities of the form
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|:={\mathbb{E}}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j},$
where $(a_{i,j})$ is a deterministic matrix and $(\varepsilon_{i,j})_{i,j\geq 1}$ is the double indexed sequence of i.i.d. symmetric $\pm 1$ r.v’s.
Since operator norm is bigger than length of every column and row we get
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|\sim({\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|^{2})^{1/2}\geq\max\Bigl{\{}\max_{i}\|(a_{i,j})_{j}\|_{2},\max_{j}\|(a_{i,j})_{i}\|_{2}\Bigr{\}}.$
For two nonnegative functions $f$ and $g$ we write $f\gtrsim$ g (or $g\lesssim f$) if there exists an absolute constant $C$ such that $Cf\geq g$; the notation $f\sim g$ means that $f\gtrsim g$ and $g\gtrsim f$. Seginer [16] proved that for $n\geq 2$,
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\log^{1/4}n\Bigr{(}\max_{i}\|(a_{i,j})_{j}\|_{2}+\max_{j}\|(a_{i,j})_{i}\|_{2}\Bigl{)}$
and constructed an example showing that in general the constant $\log^{1/4}n$ cannot be improved.
Let $g_{i,j}$ be independent $\mathcal{N}(0,1)$ random variables. We have
(1.1) ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|\lesssim{\mathbb{E}}\|(a_{i,j}g_{i,j})\|\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+\max_{j}\|(a_{i,j})_{i}\|_{2}+\sqrt{\log n}\max_{i,j}|a_{i,j}|,$
where the last bound was showned by Bandeira and van Handel [5]. The bound on the Gaussian matrices is sharp for $\{0,1\}$-weights; however, even in this case, estimate (1.1) often yields suboptimal bounds (see Section 2 below).
In [11, Theorem 1.1] it was shown that for any matrix $(a_{ij})$,
(1.2) ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|$ $\gtrsim\max_{1\leq i\leq n}\|(a_{i,j})_{j}\|_{2}+\max_{1\leq j\leq n}\|(a_{i,j})_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}.$
Here and in the sequel $\mathrm{Log}\ x=\log(x\vee e)$ and $\|S\|_{p}=({\mathbb{E}}|S|^{p})^{1/p}$ denotes $L_{p}$-norm of a r.v. $S$.
It was also conjectured that bound (1.2) may be reversed, i.e., for any scalar matrix $(a_{i,j})_{i,j\leq n}$,
(1.3) ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|$ $\lesssim\max_{1\leq i\leq n}\|(a_{i,j})_{j}\|_{2}+\max_{1\leq j\leq n}\|(a_{i,j})_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}.$
The proof of [11, Remark 4.5], based on the permutation method from [12], shows that in order to establish (1.3) it is enough to show that for any submatrix $(b_{i,j})_{i,j\leq m}$ of $(a_{i,j})_{i,j\leq n}$ one has
(1.4) ${\mathbb{E}}\|(b_{i,j}\varepsilon_{i,j})_{i,j\leq m}\|\lesssim\max_{1\leq i\leq m}\|(b_{i,j})_{j}\|_{2}+\max_{1\leq j\leq m}\|(b_{i,j})_{i}\|_{2}+R_{B}(\mathrm{Log}\,m),$
where for a matrix $A=(a_{i,j})$ and $p\geq 1$ we put
$R_{A}(p):=\sup_{\|s\|_{2}\leq 1,\|t\|_{2}\leq 1}\Big{\|}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{\|}_{p}.$
Our first result states that this conjectured bounds holds for $\{0,1\}$-matrices.
Theorem 1.1.
Inequality (1.4) holds if $b_{i,j}\in\{0,1\}$ for any $i,j$. As a consequence, for any $E\subset[n]\times[n]$,
(1.5) ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|$ $\sim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}\mathbbm{1}_{E}(i,j)\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}.$
Inequality (1.4) for $\{0,1\}$-weights is a consequence of the more general Theorem 1.6 below, applied to the symmetric $2m\times 2m$ {0,1}-matrix $A=\left(\begin{array}[]{cc}0&A\\ A^{T}&0\end{array}\right)$. Estimate (1.5) follows from (1.4) as in the proof of [11, Remark 4.5].
Remark 1.2.
In our results we do not assume the symmetry of $(\varepsilon_{i,j})_{i,j}$ even if we consider symmetric matrices $(a_{i,j})$. However analogous upper bounds holds for ${\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j})_{i,j}\|$, where $(\tilde{\varepsilon}_{i,j})_{i,j}$ is the symmetric Rademacher matrix (i.e., $\tilde{\varepsilon}_{i,j}=\tilde{\varepsilon}_{j,i}=\varepsilon_{i,j}$ for $i\geq j$), since
${\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j})_{i,j}\|$ $\leq{\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j}\mathbbm{1}_{\{i\leq j\}})_{i,j}\|+{\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j}\mathbbm{1}_{\{i>j\}})_{i,j}\|$
$={\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j}\mathbbm{1}_{\{i\leq j\}})_{i,j}\|+{\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j}\mathbbm{1}_{\{i>j\}})_{i,j}\|\leq 2{\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|.$
Moreover, the lower bound (1.2) remains valid if we replace $\varepsilon_{i,j}$ with $\tilde{\varepsilon}_{i,j}$.
Quantity $R_{A}(p)$ involves random variables $\varepsilon_{i,j}$. It may be expressed in terms of $a_{i,j}$ using two-sided bounds for $L_{p}$-norms of Rademacher sums (derived in [10] on the base of tail bounds [15])
$\Bigg{\|}\sum_{k=1}^{n}a_{k}\varepsilon_{k}\Bigg{\|}_{p}\sim\sum_{k\leq p}a_{k}^{*}+\sqrt{p}\Bigg{(}\sum_{k>p}(a_{k}^{*})^{2}\Bigg{)}^{1/2}\sim\sup\Bigg{\{}\sum_{k=1}^{n}a_{k}b_{k}\colon\|b\|_{\infty}\leq 1,\|b\|_{2}\leq\sqrt{p}\Bigg{\}},$
where $(a_{k}^{*})_{k=1}^{n}$ denotes the nonincreasing rearrangement of $(|a_{k}|)_{k=1}^{n}$. It is still unclear how to apply the above bounds to get a simple two-sided estimate for $R_{A}(p)$. However, in [11, Proposition 1.4], such an estimate was established for $\{0,1\}$-matrices:
$\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j}\mathbbm{1}_{E}(i,j)\varepsilon_{ij}s_{i}t_{j}\right\|_{p}\sim\max_{F\subset E,|F|\leq p}\|(\mathbbm{1}_{\{(i,j)\in F\}})\|.$
Hence Theorem 1.1 implies the following corollary – its first part provides a positive answer to the question posed by Ramon van Handel (private communication).
Corollary 1.3.
For any $E\subset[n]\times[n]$,
(1.6) ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|$ $\lesssim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
$\phantom{aa}+\sup_{F\subset E,|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F\}})_{i,j}\|.$
and
(1.7) ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|$ $\sim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
$\phantom{aa}+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\max_{F\subset E,|F|\leq\mathrm{Log}\,k}\|(\mathbbm{1}_{\{(i,j)\in F,i,j\notin I\}})_{i,j}\|.$
Example 2.7 below shows that one cannot reverse estimate (1.6) and a more involved form (1.7) of the two-sided estimate is necessary.
Remark 1.4.
Two-sided bounds on moments of norms of Rademacher vectors [9] give that for every $p\geq 1$,
$\Bigl{(}{\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|^{p}\Bigr{)}^{1/p}\sim{\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|+R_{A}(p).$
Thus, estimate (1.6) might be equivalently stated as
(1.8) $({\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|^{2\lfloor\mathrm{Log}\ n\rfloor})^{1/2\lfloor\mathrm{Log}n\rfloor}$ $\sim\max_{1\leq i\leq n}\|(\mathbbm{1}_{E}(i,j))_{j}\|_{2}+\max_{1\leq j\leq n}\|(\mathbbm{1}_{E}(i,j))_{i}\|_{2}$
$\phantom{aa}+\max_{F\subset E,|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F\}})_{i,j}\|.$
It is quite tempting to show (1.8) for symmetric sets $E$ via a combinatorial method, since for $n\times n$ symmetric matrix $A$ and $k=\lfloor\mathrm{Log}\,n\rfloor$, $\|A\|\sim(\mathrm{tr}(A^{2k}))^{1/{2k}}$. Such an approach worked for Gaussian matrices [5], but we were not able to apply it in the Rademacher case.
Remark 1.5.
Signed adjacency matrices were studied in [7] in connection with 2-lifts of graphs. [7, Lemma 3.1] shows that to each signed adjacency matrix of a graph $G$ one may associate the 2-lift of $G$ with the set of eigenvalues being the union of the eigenvalues of $G$ and of the signed matrix. Hence Theorem 1.1 provides an average uniform bound on new eigenvalues of random 2-lifts.
To state results for general matrices we need to introduce some additional notation. We associate to a symmetric matrix $(a_{i,j})_{i,j\leq n}$ a graph $G_{A}=([n],E_{A})$, where $(i,j)\in E_{A}$ iff $i\neq j$ and $a_{i,j}\neq 0$. By $d_{A}$ we denote the maximal degree of vertices in $G_{A}$. Observe that in the case of $\{0,1\}$-matrices $\sqrt{d_{A}}\|(a_{i,j})\|_{\infty}=\sqrt{d_{A}}=\max_{i}\|(a_{i,j})_{j}\|_{2}$.
Theorem 1.6.
For any symmetric matrix $(a_{i,j})_{i,j\leq n}$,
(1.9) ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n)+d_{A}^{19/40}\|(a_{i,j})\|_{\infty}.$
Remark 1.7.
Since $\|(a_{i,i}\varepsilon_{i,i})\|=\max_{i}|a_{i,i}|$ we may only consider matrices with zero diagonal. Moreover, for any unit vectors $s,t$ we have
$\Big{|}\sum_{i\neq j}a_{i,j}\varepsilon_{ij}s_{i}t_{j}\Big{|}$ $\leq\|(a_{i,j})\|_{\infty}\sum_{i,j}\mathbbm{1}_{\{(i,j)\in E_{A}\}}\frac{1}{2}(s_{i}^{2}+t_{j}^{2})$
$=\frac{\|(a_{i,j})\|_{\infty}}{2}\Bigl{(}\sum_{i}s_{i}^{2}\sum_{j}\mathbbm{1}_{\{(i,j)\in E_{A}\}}+\sum_{j}t_{j}^{2}\sum_{i}\mathbbm{1}_{\{(i,j)\in E_{A}\}}\Bigl{)}$
$\leq d_{A}\|(a_{i,j})\|_{\infty}.$
Hence,
(1.10) ${\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{i\neq j\}}\varepsilon_{i,j})_{i,j}\|\leq d_{A}\|(a_{i,j})\|_{\infty}$
and it is enough to consider only the case $n\geq d_{A}\geq 3$.
The proof of Theorem 1.6 takes the most part of the paper. Here we briefly sketch the main ideas of this proof. Bernoulli conjecture, formulated by Talagrand and proven in [6], states that to estimate a supremum of the Bernoulli process one needs to decompose the index set into two parts and estimate supremum over the first part using the uniform bound and over the second part by the supremum of the Gaussian process. Unfortunately, there is no algorithmic method for making such a decomposition – a rule of thumb is that the uniform bound works well for large coefficients and the Gaussian bound for small ones. We try to follow this informal recipe, decompose vectors $s,t\in B_{2}^{n}$ into almost "flat" parts and use the uniform bound when infinity norms of these parts are far apart. When they are of the same order we make some further technical adjustments (using properties of the graph $G_{A}$) and apply the Gaussian bound. The crucial tool used to estimate the corresponding Gaussian process is an improvement of van Handel’s bound [17], provided in Section 2.1.
We postpone the details of the proof till the end of the paper and discuss now some consequences of Theorem 1.6.
Theorem 1.8.
For any symmetric matrix $(a_{i,j})_{i,j\leq n}$,
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\leq\mathrm{Log}\mathrm{Log}(d_{A})\Bigr{(}\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n)\Bigl{)}.$
Proof.
Let $M:=\max_{i}\|(a_{i,j})_{j}\|_{2}$, $u_{0}=1$ and $u_{k}:=\exp(-(20/19)^{k})$ for $k=1,2,\ldots$. Let $k_{0}$ be the smallest integer such that $(\frac{20}{19})^{k_{0}}\geq\mathrm{Log}(d_{A})$. Then $k_{0}\sim\mathrm{Log}\mathrm{Log}(d_{A})$ and $u_{k_{0}}\leq d_{A}^{-1}$. We have
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|\leq{\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{|a_{i,j}|\leq u_{k_{0}}M\}}\varepsilon_{i,j})\|+\sum_{k=1}^{k_{0}}{\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{u_{k}M<|a_{i,j}|\leq u_{k-1}M\}}\varepsilon_{i,j})\|.$
For any $k$,
$d_{k}:=\max_{i}|\{j\colon\ |a_{i,j}|>u_{k}M\}|\leq u_{k}^{-2},$
so by Theorem 1.6
${\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{u_{k}M<|a_{i,j}|\leq u_{k-1}M\}}\varepsilon_{i,j})\|\lesssim M+R_{A}(\mathrm{Log}\,n)+d_{k}^{19/40}u_{k-1}M\lesssim M+R_{A}(\mathrm{Log}\,n).$
Moreover, using again Theorem 1.6
${\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{|a_{i,j}|\leq u_{k_{0}}M\}}\varepsilon_{i,j})\|\lesssim M+R_{A}(\mathrm{Log}\,n)+d_{A}^{19/40}u_{k_{0}}M\lesssim M+R_{A}(\mathrm{Log}\,n)\qed$
Obviously, $d_{A}\leq n$, so Theorem 1.8 (together with the standard symmetrization argument) implies that bounds (1.4) and (1.3) hold up double logarithms of $n$. However, decomposing matrix into two parts and using the Bandeira-van Handel bound one may derive conjectured upper bounds up to triple logarithms.
Theorem 1.9.
For any matrix $(a_{i,j})_{i,j\leq n}$,
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\mathrm{Log}\mathrm{Log}\mathrm{Log}\,n\Bigl{(}\max_{1\leq i\leq n}\|(a_{i,j})_{j}\|_{2}+\max_{1\leq j\leq n}\|(a_{i,j})_{i}\|_{2}+R_{A}(\mathrm{Log}\,n)\Bigr{)}$
and
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\mathrm{Log}\mathrm{Log}\mathrm{Log}\,n\Biggl{(}$ $\max_{1\leq i\leq n}\|(a_{i,j})_{j}\|_{2}+\max_{1\leq j\leq n}\|(a_{i,j})_{i}\|_{2}$
$+\max_{1\leq k\leq n}\min_{I\subset[n],|I|\leq k}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\notin I}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\right\|_{\mathrm{Log}\,k}\Biggr{)}.$
Proof.
Assume first that the matrix $(a_{i,j})$ is symmetric. Put $M:=\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}$. Estimate (1.1) yields
${\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{|a_{i,j}|\leq M\mathrm{Log}^{-1/2}\,n\}}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}.$
We have
$\max_{i}|\{j\colon\ |a_{i,j}|>M\mathrm{Log}^{-1/2}\,n\}|\leq\mathrm{Log}\,n,$
hence Theorem 1.8, applied to a matrix $(a_{i,j}\mathbbm{1}_{\{|a_{i,j}|>M\mathrm{Log}^{-1/2}\,n\}})_{i,j\leq n}$ implies
${\mathbb{E}}\|(a_{i,j}\mathbbm{1}_{\{|a_{i,j}|>M\mathrm{Log}^{-1/2}\,n\}}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\mathrm{Log}\mathrm{Log}\mathrm{Log}\,n\Bigl{(}\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n)\Bigr{)}.$
Therefore, for any symmetric matrix $(a_{i,j})$,
(1.11) ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\mathrm{Log}\mathrm{Log}\mathrm{Log}\,n\Bigl{(}\max_{1\leq i\leq m}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n)\Bigr{)}.$
Now, supppose that matrix $(a_{i,j})$ is arbitrary. Applying (1.11) to the symmetric $2n\times 2n$ matrix $\left(\begin{array}[]{cc}0&A\\ A^{T}&0\end{array}\right)$ we get the first part of the assertion.
The second part follows from the first one as in the proof of [11, Remark 4.5]. ∎
Organization of the paper. In the next section we present several examples of applications of the main results. In Section 3 we discuss basic tools used in the sequel, including an improvement of the van Handel bound for norms of Gaussian matrices from [17]. In Section 4 we derive a weaker version of Theorem 1.8 with $\log(d_{A})$ instead of $\log\log(d_{A})$ factors. The last section is devoted to the proof of Theorem 1.6.
2. Examples
In the first examples we discuss how to apply Corollary 1.3 to estimate ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|$ for various classes of graphs $G=([n],E)$. Due to Remark 1.2 the presented bounds are valid also for symmetric random matrices ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\tilde{\varepsilon}_{i,j})_{i,j}\|$.
Example 2.1.
Let $G=([n],E)$ be a graph with maximal degree $d$. Then
(2.1) ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|\leq d$
and
(2.2) ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|\lesssim\sqrt{d}+\sqrt{\mathrm{Log}\,n}.$
The first bound follows by (1.10). Estimate (2.2) follows from (1.1), one may also use (1.6) and bound the operator norm of $\mathbbm{1}_{F}$ by the Hilbert-Schmidt norm.
Example 2.2.
Let $G=([n],E)$ be the disjoint union of $m$ complete graphs of size $d+1$ (i.e. $n=m(d+1)$ and $\mathbbm{1}_{E}$ is a block-diagonal matrix with $m$ $(d+1)\times(d+1)$ blocks with ones outside diagonal). Then
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|\sim\sqrt{d}+\min\{d,\sqrt{\mathrm{Log}\,n}\}\sim\begin{cases}d&\mbox{if }d\leq\sqrt{\mathrm{Log}\,n},\\ \sqrt{\mathrm{Log}\,n}&\mbox{if }\sqrt{\mathrm{Log}\,n}\leq d\leq\mathrm{Log}\,n,\\ \sqrt{d}&\mbox{if }d\geq\mathrm{Log}\,n.\end{cases}$
Indeed, $G$ has degree $d$ and
$\max_{F\subset E,|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F\}})_{i,j}\|=\max_{F\subset[d+1]\times[d+1],|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F,i\neq j\}})_{i,j}\|\sim\min\{d,\sqrt{\mathrm{Log}\,n}\}.$
hence the upper bound for ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|$ follows by (1.6). To derive the lower bound it is enough to consider only the case $\mathrm{Log}\,n\geq d$. We then apply estimate (1.7) and observe that
$\min_{|I|\leq m-1}\max_{F\subset E,|F|\leq\mathrm{Log}(m-1)}\|(\mathbbm{1}_{\{(i,j)\in F,i,j\notin I\}})_{i,j}\|$ $=\max_{F\subset[d]\times[d],|F|\leq\mathrm{Log}(m-1)}\|(\mathbbm{1}_{\{(i,j)\in F,i\neq j\}})_{i,j}\|$
$\sim\min\{d,\sqrt{\mathrm{Log}\,m}\}\sim\min\{d,\sqrt{\mathrm{Log}\,n}\}.$
Example 2.3.
Let $G=([n],E)$ have maximal degree $d$ and girth at least $\varepsilon\mathrm{Log}\mathrm{Log}\,n$. Then ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|\sim_{\varepsilon}\sqrt{d}$.
By Corollary (1.3) it is enough to show that any subgraph $H=(V,F)$ of $([n],E)$ with at most $\mathrm{Log}\,n$ edges has spectral radius at most $C(\varepsilon)\sqrt{d}$. Subgraph $H$ has at most $2\mathrm{Log}\,n$ vertices, maximal degree at most $d$ and girth at least $\varepsilon\mathrm{Log}\mathrm{Log}\,n$. Let $k=2\lfloor\frac{\varepsilon}{2}\mathrm{Log}\mathrm{Log}\ n\rfloor$ then $\|(\mathbbm{1}_{F})\|\leq(\mathrm{tr}(\mathbbm{1}_{F}^{k}))^{1/k}$. Since $F$ does not contain cycles of lenghts at most $k$, we have
$\mathrm{tr}(\mathbbm{1}_{F}^{k})=\sum_{i_{1},\ldots,i_{k}}\mathbbm{1}_{\{(i_{1},i_{2})\in F\}}\cdots\mathbbm{1}_{\{(i_{k-1},i_{k})\in F\}}\mathbbm{1}_{\{(i_{k},i_{1})\in F\}}\leq 2\mathrm{Log}\,n\cdot N_{k,d},$
where $N_{k,d}\leq 2^{k}d^{k/2}$ is the number of closed pathes from root to itself in the $d$-regular tree. Hence we obtain the desired bound
$\|(\mathbbm{1}_{F})\|\leq(2\mathrm{Log}\,n)^{1/k}N_{k,d}^{1/k}\lesssim_{\varepsilon}\sqrt{d}.$
The next example generalizes the previous one. It is close to [14, Theorem 1.2], where there was a stronger assumption on the neighborhood diameter, but a more precise bound on the operator norm.
Example 2.4.
Let $G=([n],E)$ have maximal degree $d$ and suppose that the $r$-neighborhood of every vertex contains at most one cycle, where $r\geq\varepsilon\mathrm{Log}\mathrm{Log}\,n$. Then ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|\sim_{\varepsilon}\sqrt{d}$.
Observe that all cycles of lenght at most $2r$ do not intersect. Let us remove one edge from each cycle of length at most $2r$ and let $G_{1}:=([n],E_{1})$, where $E_{1}$ contains all removed edges and $G_{2}:=([n],E_{2})$, where $E_{2}=E\setminus E_{1}$. Then $G_{2}$ has girth at least $2r$ and the maximal degree of $G_{1}$ is at most $1$. Thus by (2.1) and Example 2.3,
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|\leq{\mathbb{E}}\|(\mathbbm{1}_{E_{1}}(i,j)\varepsilon_{i,j})\|+{\mathbb{E}}\|(\mathbbm{1}_{E_{2}}(i,j)\varepsilon_{i,j})\|\lesssim_{\varepsilon}1+\sqrt{d}\lesssim\sqrt{d}.$
The next example is similar to [14, Theorem 1.16], where it was showned that for a random $d$-regular graph $G=([n],E)$, $\|(\mathbbm{1}_{E}(i,j)\tilde{\varepsilon}_{i,j})\|\leq 2\sqrt{d-1}+\varepsilon$ with probability $1-o_{n}(1)$.
Example 2.5.
Let $G=([n],E)$ be a random $d$-regular graph. Then ${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|\sim\sqrt{d}$.
By (2.2) we may assume that $\mathrm{Log}\,n\geq d$ and by (2.1) that $d\geq 3$. Following [8] we say that a graph is $r$-tangle free if the $r$-neighborhood of every vertex contains at most one cycle By Example 2.4 and (2.1) we have
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})\|\leq C\sqrt{d}\,\mathbb{P}(G\mbox{ is $\mathrm{Log}\mathrm{Log}n$-tangle free})+d\,\mathbb{P}(G\mbox{ is not $\mathrm{Log}\mathrm{Log}n$-tangle free}).$
By [8, Lemma 27]
$\mathbb{P}(\mbox{$G$ is not $\mathrm{Log}\mathrm{Log}n$-tangle free})\lesssim\frac{(d-1)^{\mathrm{Log}\mathrm{Log}n}}{n}\lesssim d^{-1/2},$
where the last bound follows since we assume that $\mathrm{Log}\,n\geq d$.
The next example essentially recovers [2, Corollary 1.3].
Example 2.6.
Let $A$ be the adjacency matrix of a $d$-regular graph and suppose that all eigenvalues of $A$ besides the largest one are bounded in absolute value by $\lambda$. Then ${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|\lesssim\lambda$.
We have $\lambda\gtrsim\sqrt{d}$ by the Alon-Boppana theorem, so by (2.2) we may assume that $\mathrm{Log}\,n\geq d$. By (1.6) it is enough to show that any subgraph $H=(V,F)$ of $([n],E)$ with at most $\mathrm{Log}\,n$ edges has spectral radius at most $C\lambda$. Subgraph $H$ has at most $2\mathrm{Log}\,n$ vertices. Let $w=n^{-1/2}(1,\ldots,1)$ be the eigenvector of $A$ corresponding to the largest eingenvalue $d$. Any $v\in{\mathbb{R}}^{V}$ with $\|v\|_{2}=1$ we may represent as $v=\langle v,w\rangle w+v^{\prime}$, where $\langle v^{\prime},w\rangle=0$ and $\|v^{\prime}\|_{2}\leq 1$. Thus
$\|\mathbbm{1}_{F}v\|_{2}\leq\|\mathbbm{1}_{F}v^{\prime}\|+\langle v,w\rangle\|\mathbbm{1}_{F}w\|_{2}\leq\lambda+\|v\|_{1}n^{-1/2}|V|^{1/2}dn^{-1/2}\leq\lambda+2d\frac{\mathrm{Log}\,n}{n}\lesssim\lambda,$
where the last estimate follows since $\mathrm{Log}\,n\geq d$ and $\lambda\gtrsim 1$.
The next example shows that estimate (1.6) cannot be reversed.
Example 2.7.
Let $1\leq d\leq n$ and $E=([d]\times[d])\cup\{(i,i)\colon\ d<i\leq n\}$ (i.e $\mathbbm{1}_{E}$ is block diagonal with one $d\times d$ block of ones and ndn-d blocks of single ones). Then
$\sup_{F\subset E,|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F\}})_{i,j}\|=\sup_{F\subset[d]\times[d],|F|\leq\mathrm{Log}\,n}\|(\mathbbm{1}_{\{(i,j)\in F\}})_{i,j}\|\sim\min\{d,\sqrt{\mathrm{Log}\,n}\}$
and the RHS of (1.6) is of the order $\sqrt{d}+\min\{d,\sqrt{\mathrm{Log}\,n}\}$, whereas
${\mathbb{E}}\|(\mathbbm{1}_{E}(i,j)\varepsilon_{i,j})_{i,j\leq n}\|={\mathbb{E}}\|(\varepsilon_{i,j})_{i,j\leq d}\|\sim\sqrt{d}.$
The last example concerns randomized circulant matrices, investigated in [11].
Example 2.8.
Suppose that $(a_{i,j})$ is a circulant matrix, i.e. $a_{i,j}=b_{i-j\,\mathrm{mod}\,n}$ for a deterministic sequence $(b_{i})_{i=0}^{n-1}$. Then for any $i$ and $j$, $\|(a_{i,j})_{j}\|_{2}=\|(a_{i,j})_{i}\|_{2}=\|(b_{i})_{i}\|_{2}$. Moreover, as was shown in the proof of [11, Theorem 1.3]
$\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j}a_{ij}\varepsilon_{ij}s_{i}t_{j}\right\|_{\mathrm{Log}\,n}\lesssim\inf_{I\subset[n],|I|\leq n/4}\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\left\|\sum_{i,j\not\in I}a_{ij}\varepsilon_{ij}s_{i}t_{j}\right\|_{\mathrm{Log}(n/4)},$
Thus Theorem 1.9 improves [11, Theorem 1.3] and yields that for circulant matrices
$\|(b_{i})\|_{2}+R_{A}(\mathrm{Log}\,n)\lesssim{\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})\|\lesssim\mathrm{Log}\mathrm{Log}\mathrm{Log}\,n\big{(}\|(b_{i})\|_{2}+R_{A}(\mathrm{Log}\,n)\big{)}.$
3. Tools
We will use the following estimate for suprema of Rademachers. It is a special case of [1, Lemma 5.10].
Proposition 3.1.
Let $T_{1},\ldots,T_{n}$ be nonempty bounded subsets of ${\mathbb{R}}^{N}$. Then
${\mathbb{E}}\max_{k\leq n}\sup_{t\in T_{k}}\sum_{i=1}^{N}t_{i}\varepsilon_{i}\lesssim\max_{k\leq n}{\mathbb{E}}\sup_{t\in T_{k}}\sum_{i=1}^{N}t_{i}\varepsilon_{i}+\max_{k\leq n}\sup_{t\in T_{k}}\Big{\|}\sum_{i=1}^{N}t_{i}\varepsilon_{i}\Big{\|}_{\mathrm{Log}\,n}.$
Another useful result is the estimate on the number of connected subsets of a graph.
Lemma 3.2.
Let $H=(V_{H},E_{H})$ be a graph with $n_{H}$ vertices and maximal degree $d_{H}$.
i) For a fixed $v\in V$ the number of connected subsets $I\subset V_{H}$ with cardinality $k$ containing $v$ is at most $(4d_{H})^{k-1}$.
ii) The number of all connected subsets $I\subset V_{H}$ with cardinality $k$ is not bigger than $n_{H}(4d_{H})^{k-1}$.
Proof.
i) We choose a connected subset $I\ni v$ by constructing its spanning tree, rooted at $v$. In order to do it we first choose an unlabelled rooted tree with $k$ vertices and then label its vertices by elements of $V_{H}$. The number of unlabelled rooted trees is less than the number of ordered trees with $k$ vertices, i.e., less than the (k1)(k-1)-th Catalan number $C_{k-1}\leq 4^{k-1}$. The root of the tree has label $v$ and the rest of vertices may be labelled in at most $d_{H}^{k-1}$ ways.
Part i) of the assertion immediately yields part ii). ∎
3.1. Van Handel-type bound
In this part we will establish the following improvement on van Handel’s bound [17].
Proposition $3.$
For any $n\times m$ matrix $(a_{i,j})_{i\leq m,j\leq n}$ and $b\in(0,1]$ we have
${\mathbb{E}}\sup_{s\in B_{2}^{m}\cap bB_{\infty}^{m}}\sup_{t\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i\leq m,j\leq n}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$ $\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+\max_{j}\|(a_{i,j})_{i}\|_{2}$
$\phantom{\lesssim}+\mathrm{Log}((n+m)b^{2})\|(a_{i,j})_{i,j}\|_{\infty}.$
Let us first formulate and prove a symmetric variant of Propostion $3.$
Proposition 3.4.
Let $(\tilde{\varepsilon}_{i,j})_{i,j}$ be a symmetric Rademacher matrix. Then for any symmetric matrix $(a_{i,j})_{i,j\leq n}$ and any $b\in(0,1]$,
${\mathbb{E}}\sup_{s,t\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i,j\leq n}a_{i,j}\tilde{\varepsilon}_{i,j}s_{i}t_{j}\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+\mathrm{Log}(nb^{2})\|(a_{i,j})_{i,j}\|_{\infty}.$
The proof uses the following, quite standard, technical lemma.
Lemma 3.5.
Let $Y_{1},\ldots,Y_{n}$ be r.v’s and $m_{i},\sigma_{i}\geq 0$ be such that
$\mathbb{P}(|Y_{i}|\geq m_{i}+u\sigma_{i})\leq e^{-u^{2}/2}\quad\mbox{ for every }u\geq 0\text{ and }i=1,\ldots,n.$
Then
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}^{2}Y_{i}\lesssim\max_{i}m_{i}+\sqrt{\mathrm{Log}(nb^{2})}\max_{i}\sigma_{i}$
and
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sqrt{\sum_{i=1}^{n}s_{i}^{2}Y_{i}^{2}}\lesssim\max_{i}m_{i}+\sqrt{\mathrm{Log}(nb^{2})}\max_{i}\sigma_{i}.$
Proof.
Let $(x_{1}^{*},\ldots,x_{n}^{*})$ denote a nondecreasing rearrangement of $|x_{1}|,\ldots,|x_{n}|$. We set $k=n$ if $b^{2}\leq 1/n$, otherwise we choose $1\leq k\leq n-1$ such that $\frac{1}{k+1}<b^{2}\leq\frac{1}{k}$. Then $\mathrm{Log}(nb^{2})\sim\mathrm{Log}(n/k)$ and for any $s\in B_{2}^{n}\cap bB_{\infty}^{n}$,
$\sum_{i=1}^{n}s_{i}^{2}Y_{i}$ $\leq\sum_{i=1}^{n}|s_{i}^{*}|^{2}Y_{i}^{*}\leq\sum_{i=1}^{k}|s_{i}^{*}|^{2}Y_{i}^{*}+\Bigl{(}1-\sum_{i=1}^{k}|s_{i}^{*}|^{2}\Bigr{)}Y_{k}^{*}$
$=\sum_{i=1}^{k}\Bigl{(}|s_{i}^{*}|^{2}Y_{i}^{*}+\Bigl{(}\frac{1}{k}-|s_{i}^{*}|^{2}\Bigr{)}Y_{k}^{*}\Bigr{)}\leq\frac{1}{k}(Y_{1}^{*}+\ldots+Y_{k}^{*}).$
For any $1\leq l\leq n$ and $u\geq 0$,
$\mathbb{P}(Y_{l}^{*}\geq\max_{i}m_{i}+u\max_{i}\sigma_{i})\leq\frac{1}{l}\sum_{i=1}^{n}\mathbb{P}(Y_{i}\geq\max_{i}m_{i}+u\max_{i}\sigma_{i})\leq\frac{n}{l}e^{-u^{2}/2}.$
Hence integration by parts yields ${\mathbb{E}}Y_{l}^{*}\leq({\mathbb{E}}|Y_{l}^{*}|^{2})^{1/2}\lesssim\max_{i}m_{i}+\mathrm{Log}^{1/2}(n/l)\max_{i}\sigma_{i}$. Thus
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}^{2}Y_{i}$ $\leq\frac{1}{k}\sum_{l=1}^{k}{\mathbb{E}}Y_{l}^{*}\lesssim\max_{i}m_{i}+\frac{1}{k}\sum_{l=1}^{k}\sqrt{\mathrm{Log}\bigl{(}\frac{n}{l}\bigr{)}}\max_{i}\sigma_{i}$
$\lesssim\max_{i}m_{i}+\sqrt{\mathrm{Log}\bigl{(}\frac{n}{k}\bigr{)}}\max_{i}\sigma_{i}.$
Similarly,
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sqrt{\sum_{i=1}^{n}s_{i}^{2}Y_{i}^{2}}$ $\leq{\mathbb{E}}\sqrt{\frac{1}{k}(|Y_{1}^{*}|^{2}+\ldots+|Y_{k}^{*}|^{2})}\leq\sqrt{\frac{1}{k}\sum_{l=1}^{k}{\mathbb{E}}|Y_{l}^{*}|^{2}}$
$\lesssim\max_{i}m_{i}+\sqrt{\frac{1}{k}\sum_{l=1}^{k}\mathrm{Log}\bigl{(}\frac{n}{l}\bigr{)}}\max_{i}\sigma_{i}$
$\lesssim\max_{i}m_{i}+\sqrt{\mathrm{Log}\bigl{(}\frac{n}{k}\bigr{)}}\max_{i}\sigma_{i}.\qed$
Proof of Proposition 3.4.
Let $(g_{i,j})_{i,j\leq n}$ be a symmetric Gaussian matrix (i.e., $g_{i,j}=g_{j,i}$ and $(g_{i,j})_{i\geq j}$ are iid $\mathcal{N}(0,1)$ r.v’s), independent of $\tilde{\varepsilon}_{i,j}$. We have for any matrix norm $\|\cdot\|$,
${\mathbb{E}}\|(a_{i,j}g_{i,j})\|={\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j}|g_{i,j}|)\|\geq{\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j}{\mathbb{E}}|g_{i,j}|)\|=\sqrt{\frac{2}{\pi}}{\mathbb{E}}\|(a_{i,j}\tilde{\varepsilon}_{i,j})\|.$
For any symmetric matrix $B$ we have $\langle Bs,t\rangle=\frac{1}{4}\bigl{(}\langle B(s+t),s+t\rangle-\langle B(s-t),s-t\rangle\bigr{)}$, hence,
$\sup_{s,t\in B_{2}^{n}\cap bB_{\infty}^{n}}\langle Bs,t\rangle\leq 2\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}|\langle Bs,s\rangle|.$
Therefore,
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i,j\leq n}a_{i,j}\tilde{\varepsilon}_{i,j}s_{i}t_{j}$ $\lesssim{\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\Bigl{|}\sum_{i,j\leq n}a_{i,j}g_{i,j}s_{i}s_{j}\Bigr{|}$
$\leq{\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i,j\leq n}a_{i,j}g_{i,j}s_{i}s_{j}+{\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i,j\leq n}(-a_{i,j}g_{i,j}s_{i}s_{j})$
$=2{\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i,j\leq n}a_{i,j}g_{i,j}s_{i}s_{j}.$
Now we follow van Handel’s approach from [17]. Let $g_{1},g_{2},\ldots,g_{n}$ be iid $\mathcal{N}(0,1)$ r.v’s and $Y=(Y_{1},\ldots,Y_{n})\sim\mathcal{N}(0,B_{-})$, where $B_{-}$ is the negative part of $B=(a_{i,j}^{2})$. Define the new Gaussian process $Z_{s}$ by
$Z_{s}=2\sum_{i=1}^{n}s_{i}g_{i}\sqrt{\sum_{j=1}^{n}a_{ij}^{2}s_{j}^{2}}+\sum_{i=1}^{n}s_{i}^{2}Y_{i}.$
It is shown in [17] (see the proof of Theorem 4.1 therein) that for any $s,s^{\prime}\in{\mathbb{R}}^{n}$
${\mathbb{E}}\Bigl{|}\sum_{i,j\leq n}a_{i,j}g_{i,j}(s_{i}s_{j}-s_{i}^{\prime}s_{j}^{\prime})\Bigr{|}^{2}\leq{\mathbb{E}}|Z_{s}-Z_{s^{\prime}}|^{2}.$
Hence the Slepian-Fernique inequality [13, Theorem 3.15]. yields
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i,j\leq n}a_{i,j}g_{i,j}s_{i}s_{j}\leq{\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}Z_{s}.$
Variables $Y_{i}$ are centered Gaussian and (see the proof of Corollary 4.2 in [17]) $({\mathbb{E}}Y_{i}^{2})^{1/2}\leq\|(a_{i,j})_{j}\|_{4}$. Hence Lemma 3.5 applied with $m_{i}=0$ and $\sigma_{i}=\|(a_{i,j})_{j}\|_{4}$ yields
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}^{2}Y_{i}$ $\lesssim\sqrt{\mathrm{Log}(nb^{2})}\max_{i}\|(a_{i,j})_{j}\|_{4}$
$\leq\sqrt{\mathrm{Log}(nb^{2})}\max_{i,j}|a_{i,j}|^{1/2}\max_{j}\|(a_{i,j})_{j}\|_{2}^{1/2}$
$\leq\max_{i}\|(a_{i,j})_{j}\|_{2}+\mathrm{Log}(nb^{2})\|(a_{i,j})_{i,j}\|_{\infty}.$
We have
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}g_{i}\sqrt{\sum_{j}a_{ij}^{2}s_{j}^{2}}\leq{\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sqrt{\sum_{i,j}a_{ij}^{2}s_{j}^{2}g_{i}^{2}}={\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sqrt{\sum_{j}s_{j}^{2}V_{j}^{2}},$
where $V_{j}=\sqrt{\sum_{i}a_{ij}^{2}g_{i}^{2}}$. The Gaussian concentration [13, Lemma 3.1] yields
$\mathbb{P}(|V_{j}|\geq\|(a_{i,j})_{i}\|_{2}+t\|(a_{i,j})_{i}\|_{\infty})\leq e^{-t^{2}/2},$
so Lemma 3.5 applied with $Y_{j}=V_{j}$, $m_{j}=\|(a_{i,j})_{i}\|_{2}$ and $\sigma_{j}=\|(a_{i,j})_{i}\|_{\infty}$ yields
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i=1}^{n}s_{i}\sqrt{\sum_{j}a_{ij}^{2}s_{j}^{2}}g_{i}\lesssim\max_{j}\|(a_{i,j})_{i}\|_{2}+\sqrt{\mathrm{Log}(nb^{2})}\|(a_{i,j})_{i,j}\|_{\infty}.\qed$
Proof of Proposition $3.$
We apply Proposition 3.4 to the symmetric $(n+m)\times(n+m)$ matrix $\tilde{A}=(\tilde{a}_{i,j})$ of the form $\tilde{A}=\left(\begin{array}[]{cc}0&A\\ A^{T}&0\end{array}\right)$. Observe that $\max_{i,j}|\tilde{a}_{i,j}|=\max_{i,j}|a_{i,j}|$,
$\max_{i}\|(\tilde{a}_{i,j})_{j}\|_{2}=\max\Bigl{\{}\max_{i}\|(a_{i,j})_{j}\|_{2},\max_{j}\|(a_{i,j})_{i}\|_{2}\Bigr{\}}$
and
${\mathbb{E}}\sup_{s,t\in B_{2}^{n+m}\cap bB_{\infty}^{n+m}}$ $\sum_{i,j\leq n+m}\tilde{a}_{i,j}\tilde{\varepsilon}_{i,j}s_{i}t_{j}\geq{\mathbb{E}}\sup_{s\in B_{2}^{m}\cap bB_{\infty}^{m}}\sup_{t\in B_{2}^{n}\cap bB_{\infty}^{n}}\sum_{i\leq m,j\leq n}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.\qed$
4. Bounds up to polylog factors
In this section we derive weaker estimates than in Theorem 1.8 (with powers of $\log d_{A}$ instead of $\log\log(d_{A})$). They will be used in the proof of Theorem 1.6 to estimate the parts of Bernoulli process $(\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j})_{s,t\in B_{2}^{n}}$, where coefficients $s_{i}$ and $t_{j}$ are of the same order.
Let us first introduce the notation which will be used till the end of the paper. Recall that $G_{A}=([n],E_{A})$ is a graph associated to a given symmetric matrix $A=(a_{i,j})_{i,j\leq n}$. By $\rho=\rho_{A}$ we denote the distance on $[n]$ induced by $E_{A}$. For $r=1,2,\ldots$ we put $G_{r}=G_{r}(A)=([n],E_{A,r})$, where $(i,j)\in E_{A,r}$ iff $\rho(i,j)\leq r$. In particular $G_{1}=([n],E_{A})$ and the maximal degree of $G_{r}$ is at most $d_{A}+d_{A}(d_{A}-1)+\ldots+d_{A}(d_{A}-1)^{r-1}\leq d_{A}^{r}$. We say that a subset of $[n]$ is $r$-connected if it is connected in $G_{r}$.
We denote by $\mathcal{I}(k)=\mathcal{I}(k,n)$ the family of all subsets of $[n]$ of cardinality $k$ and by $\mathcal{I}_{r}(k)=\mathcal{I}_{r}(k,A)$ the family of all $r$-connected subsets of $[n]$ of cardinality $k$.
For a set $I\subset[n]$ and a vertex $j\in[n]$ we write $I\sim_{A}j$ if $(i,j)\in E_{A}$ for some $i\in I$. By $I^{\prime}=I^{\prime}(A)$ we denote the set of all neighbours of $I$ in $G_{1}$ and by $I^{\prime\prime}=I^{\prime\prime}(A)$ the set of all neighbours of $I^{\prime}$ in $G_{1}$, i.e.,
(4.1) $I^{\prime}=\{j\in[n]\colon\ \exists_{i\in I}\ (i,j)\in E_{A}\},\quad I^{\prime\prime}=\{i\in[n]\colon\ \exists_{i_{0}\in I,j\in[n]}\ (i_{0},j),(i,j)\in E_{A}\}.$
Observe that $I$ is a subset of $I^{\prime\prime}$, but does not have to be a subset of $I^{\prime}$. Moreover $|I^{\prime}|\leq d_{A}|I|$ and $|I^{\prime\prime}|\leq d_{A}^{2}|I|$.
By Remark 1.7 we may and will assume that $a_{i,i}=0$ for all $i$.
For $1\leq k,l\leq n$ define random variables
$X_{k,l}=X_{k,l}(A):=\frac{1}{\sqrt{kl}}\max_{I\in\mathcal{I}(k),J\in\mathcal{I}(l)}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}^{\prime}$
and their $4$-connected counteparts
$\overline{X}_{k,l}=\overline{X}_{k,l}(A):=\frac{1}{\sqrt{kl}}\max_{I\in\mathcal{I}_{4}(k),J\in\mathcal{I}_{4}(l)}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}^{\prime}.$
Set
(4.2) $X=X(A):=\max_{1\leq k,l\leq n}X_{k,l}=\max_{\emptyset\neq I,J\subset V}\frac{1}{\sqrt{|I||J|}}\max_{\eta_{i},\eta^{\prime}_{j}=\pm 1}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}.$
Observe that $X$ is nonnegative.
Variables $\overline{X}_{k,l}$ are easier to estimate than $X_{k,l}$, since the number of $4$-connected subsets is much smaller than the number of all subsets – calculations based on this idea are made in Lemma 4.2. Lemma 4.1 shows that in expectation these variables do not differ too much.
Lemma 4.1.
For any $1\leq k,l\leq n$,
${\mathbb{E}}X_{k,l}\lesssim\max_{1\leq k^{\prime}\leq k,1\leq l^{\prime}\leq l}{\mathbb{E}}\overline{X}_{k^{\prime},l^{\prime}}+R_{A}(\mathrm{Log}(kl)).$
Proof.
Let us first fix sets $I\in\mathcal{I}(k)$ and $J\in\mathcal{I}(l)$. Let $I_{1},\ldots,I_{r}$ be connected components of $I\cap J^{\prime}$ in $G_{2}$ and $J_{u}:=J\cap I_{u}^{\prime}$. Then sets $J_{1},\ldots,J_{r}$ are disjoint. They are also 4-connected subsets of $J$, since otherwise there would exists a nonempty set $V\subsetneq J_{u}$ such that $\rho_{A}(V,J_{u}\setminus V)\geq 4$ and taking for $\tilde{V}$ neighbours of $V$ in $I_{u}$ we would have $\emptyset\neq\tilde{V}\subsetneq I_{u}$ and $\rho_{A}(\tilde{V},I_{u}\setminus\tilde{V})\geq 2$, contradicting 2-connectivity of $I_{u}$. Hence, for every $\eta_{i},\eta_{j}^{\prime}=\pm 1$ we have
$\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}^{\prime}$ $=\sum_{i\in I\cap J^{\prime},j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}^{\prime}=\sum_{u=1}^{r}\sum_{i\in I_{u},j\in J_{u}}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}^{\prime}\leq\sum_{u=1}^{r}\overline{X}_{|I_{u}|,|J_{u}|}\sqrt{|I_{u}||J_{u}|}$
$\leq\max_{k^{\prime}\leq k,l^{\prime}\leq l}\overline{X}_{k^{\prime},l^{\prime}}\sum_{u=1}^{r}\sqrt{|I_{u}||J_{u}|}\leq\max_{k^{\prime}\leq k,l^{\prime}\leq l}\overline{X}_{k^{\prime},l^{\prime}}\Bigl{(}\sum_{u=1}^{r}|I_{u}|\Bigr{)}^{1/2}\Bigl{(}\sum_{u=1}^{r}|J_{u}|\Bigr{)}^{1/2}$
$\leq\max_{k^{\prime}\leq k,l^{\prime}\leq l}\overline{X}_{k^{\prime},l^{\prime}}\sqrt{|I||J|}.$
Taking the supremum over all sets $I\in\mathcal{I}(k)$, $J\in\mathcal{I}(l)$ and $\eta_{i},\eta_{j}^{\prime}=\pm 1$ we get
(4.3) $X_{k,l}\leq\max_{k^{\prime}\leq k,l^{\prime}\leq l}\overline{X}_{k^{\prime},l^{\prime}}.$
Observe that
$\max_{I\in\mathcal{I}_{4}(k^{\prime}),J\in\mathcal{I}_{4}(l^{\prime})}\max_{\eta_{i},\eta^{\prime}_{j}=\pm 1}\frac{1}{\sqrt{k^{\prime}l^{\prime}}}\Bigl{\|}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}\Bigr{\|}_{\mathrm{Log}(kl)}$ $\leq\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\Bigl{\|}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{\|}_{\mathrm{Log}(kl)}$
$=R_{A}(\mathrm{Log}(kl)).$
Thus, by Proposition 3.1
${\mathbb{E}}\max_{k^{\prime}\leq k,l^{\prime}\leq l}\overline{X}_{k^{\prime},l^{\prime}}\lesssim\max_{k^{\prime}\leq k,l^{\prime}\leq l}{\mathbb{E}}\overline{X}_{k^{\prime},l^{\prime}}+R_{A}(\mathrm{Log}(kl)).$
Lemma 4.2.
We have for any $1\leq k,l\leq n$,
${\mathbb{E}}\overline{X}_{k,l}\lesssim\sqrt{\mathrm{Log}\,d_{A}}\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n).$
Proof.
Obviously $\overline{X}_{k^{\prime},l^{\prime}}\leq\|(a_{i,j}\varepsilon_{i,j})\|$, so by (1.10) we may assume that $n\geq d_{A}\geq 3$. By the symmetry it is enough to consider only the case $l\geq k$. By Lemma 3.2, $2^{k}|\mathcal{I}_{4}(k)|\leq n(8d_{A}^{4})^{k}\leq nd_{A}^{6k}$.
We have
$\overline{X}_{k,l}\leq\frac{1}{\sqrt{k}}\max_{I\in\mathcal{I}_{4}(k)}\max_{\eta_{i}=\pm 1}\sup_{\|t\|_{2}\leq 1}\sum_{i\in I}\sum_{j}a_{i,j}\varepsilon_{i,j}\eta_{i}t_{j}.$
For any fixed $I\in\mathcal{I}_{4}(k)$ and $\eta_{i}=\pm 1$,
${\mathbb{E}}\frac{1}{\sqrt{k}}\sup_{\|t\|_{2}\leq 1}\sum_{i\in I}\sum_{j}a_{i,j}\varepsilon_{i,j}\eta_{i}t_{j}$ $=\frac{1}{\sqrt{k}}{\mathbb{E}}\Bigl{(}\sum_{j}\Bigr{(}\sum_{i\in I}a_{i,j}\varepsilon_{i,j}\eta_{i}\Bigl{)}^{2}\Bigr{)}^{1/2}$
$\leq\frac{1}{\sqrt{k}}\Bigl{(}\sum_{j}{\mathbb{E}}\Bigr{(}\sum_{i\in I}a_{i,j}\varepsilon_{i,j}\eta_{i}\Bigl{)}^{2}\Bigr{)}^{1/2}=\frac{1}{\sqrt{k}}\Bigl{(}\sum_{j}\sum_{i\in I}a_{i,j}^{2}\Bigr{)}^{1/2}$
$=\frac{1}{\sqrt{k}}\Bigl{(}\sum_{i\in I}\sum_{j}a_{i,j}^{2}\Bigr{)}^{1/2}\leq\max_{i}\|(a_{i,j})_{j}\|_{2}.$
In the case $n\geq d_{A}^{6k}$, $\log(2^{k}|\mathcal{I}_{4}(k)|)\lesssim\log n$ and
$\max_{I\in\mathcal{I}_{4}(k)}\max_{\eta_{i}=\pm 1}\sup_{\|t\|_{2}\leq 1}\frac{1}{\sqrt{k}}\Bigr{\|}\sum_{i\in I}\sum_{j}a_{i,j}\varepsilon_{i,j}\eta_{i}t_{j}\Bigl{\|}_{\mathrm{Log}(2^{k}|\mathcal{I}_{4}(k)|)}\lesssim R_{A}(\log n).$
In the case $n\leq d_{A}^{6k}$ we have $\log(2^{k}|\mathcal{I}_{4}(k)|)\lesssim k\log d_{A}$ and
$\max_{I\in\mathcal{I}_{4}(k)}$ $\max_{\eta_{i}=\pm 1}\sup_{\|t\|_{2}\leq 1}\frac{1}{\sqrt{k}}\Bigr{\|}\sum_{i\in I}\sum_{j}a_{i,j}\varepsilon_{i,j}\eta_{i}t_{j}\Bigl{\|}_{\mathrm{Log}(2^{k}|\mathcal{I}_{4}(k)|)}$
$\lesssim\max_{I\in\mathcal{I}_{4}(k)}\max_{\eta_{i}=\pm 1}\sup_{\|t\|_{2}\leq 1}\sqrt{\log d_{A}}\Bigl{(}\sum_{i\in I}\sum_{j}(a_{i,j}\eta_{i}t_{j})^{2}\Bigr{)}^{1/2}$
$=\max_{I\in\mathcal{I}_{4}(k)}\max_{j}\sqrt{\log d_{A}}\Bigl{(}\sum_{i\in I}a_{i,j}^{2}\Bigr{)}^{1/2}\leq\sqrt{\log d_{A}}\max_{j}\|(a_{i,j})_{i}\|_{2}.$
The assertion follows by Proposition 3.1. ∎
Corollary 4.3.
We have
${\mathbb{E}}X={\mathbb{E}}\max_{\emptyset\neq I,J\subset V}\frac{1}{\sqrt{|I||J|}}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}\lesssim\sqrt{\mathrm{Log}\,d_{A}}\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n).$
Proof.
Lemmas 4.1 and 4.2 imply that for a fixed $1\leq k,l\leq n$
${\mathbb{E}}X_{k,l}\lesssim\max_{k^{\prime}\leq k,l^{\prime}\leq l}{\mathbb{E}}\overline{X}_{k^{\prime},l^{\prime}}+R_{A}(\mathrm{Log}(kl))\lesssim\sqrt{\mathrm{Log}\,d_{A}}\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\mathrm{Log}\,n).$
Moreover,
$\max_{1\leq k,l\leq n}\max_{I\in\mathcal{I}(k),J\in\mathcal{I}(l)}\frac{1}{\sqrt{kl}}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}$ $\Big{\|}\sum_{i\in I,j\in J}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}\Big{\|}_{\mathrm{Log}(n^{2})}$
$\leq\sup_{\|t\|_{2}\leq 1,\|s\|_{2}\leq 1}\Big{\|}\sum_{i,j\in V}a_{i,j}\varepsilon_{i,j}t_{i}s_{j}\Big{\|}_{2\mathrm{Log}\,n}\lesssim R_{A}(\mathrm{Log}\,n)$
and the assertion follows by Proposition 3.1. ∎
Proposition 4.4.
For any symmeric matrix $(a_{i,j})_{i,j\leq n}$ we have
${\mathbb{E}}\bigl{\|}(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\bigr{\|}\lesssim\mathrm{Log}^{3/2}(d_{A})\max_{i}\|(a_{i,j})_{j}\|_{2}+\mathrm{Log}(d_{A})R_{A}(\mathrm{Log}\,n).$
Proof.
By Remark 1.7 we may assume that $a_{i,i}=0$ for all $i$ and $n\geq d_{A}\geq 3$.
For vectors $s,t$ and integers $k,l$ we define sets
$I_{k}(s)=\{i\leq n\colon\ e^{-k-1}<|s_{i}|\leq e^{-k}\},\quad J_{l}(t)=\{j\leq n\colon\ e^{-l-1}<|t_{j}|\leq e^{-l}\}.$
Observe that for any $s,t,k,l$,
$\sum_{i\in I_{k}(s),j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\leq e^{-k-l}\max_{\eta_{i},\eta^{\prime}_{j}=\pm 1}\sum_{i\in I_{k}(s),j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j},$
therefore
$\bigl{\|}(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\bigr{\|}\leq\sup_{\|s\|_{2},\|t\|_{2}\leq 1}\sum_{k,l}e^{-k-l}\sup_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I_{k}(s),j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}.$
We have
$\sum_{k}\sum_{l\geq k+\log d_{A}}$ $e^{-k-l}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I_{k}(s),j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}$
$\leq\sum_{k}\sum_{l\geq k+\log d_{A}}e^{-k}\sum_{i\in I_{k}(s),j\in J_{l}(t)}|a_{i,j}|e^{-l}$
$=\sum_{k}e^{-k}\sum_{i\in I_{k}(s)}\sum_{j}|a_{i,j}|\sum_{l\geq k+\log d_{A}}e^{-l}\mathbbm{1}_{\{j\in J_{l}(t)\}}$
$\leq\sum_{k}\sum_{i\in I_{k}(s)}\sum_{j}|a_{i,j}|e^{-2k-\log d_{A}}\leq\|(a_{i,j})\|_{\infty}\sum_{k}\sum_{i\in I_{k}(s)}e^{-2k}$
$\leq\|(a_{i,j})\|_{\infty}\sum_{k}\sum_{i\in I_{k}(s)}e^{2}s_{i}^{2}=e^{2}\|s\|_{2}^{2}\|(a_{i,j})\|_{\infty}.$
In the same way we show that
$\sum_{l}\sum_{k\geq l+\log d_{A}}e^{-k-l}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I_{k}(s),j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}\leq e^{2}\|t\|_{2}^{2}\|(a_{i,j})\|_{\infty}.$
Moreover, for any $s,t$,
$\sum_{k,l\colon|k-l|<\log d_{A}}e^{-k-l}\sup_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{i\in I_{k}(s),j\in J_{l}(t)}$ $a_{i,j}\varepsilon_{i,j}\eta_{i}\eta^{\prime}_{j}$
$\leq X\sum_{k,l\colon|k-l|<\log d_{A}}^{\infty}e^{-k-l}\sqrt{|I_{k}(s)||J_{l}(t)|}.$
For any fixed integer $r$
$\sum_{k}e^{-k-(k+r)}\sqrt{|I_{k}(s)||J_{k+r}(t)|}$ $\leq\Bigl{(}\sum_{k}e^{-2k}|I_{k}(s)|\Bigr{)}^{1/2}\Bigl{(}\sum_{k}e^{-2(k+r)}|J_{k+r}(t)|\Bigr{)}^{1/2}$
$\leq e^{2}\|t\|_{2}\|s\|_{2}.$
Hence,
$\bigl{\|}(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\bigr{\|}$ $\leq\sup_{\|s\|_{2},\|t\|_{2}\leq 1}e^{2}((\|s\|^{2}+\|t\|^{2})\|(a_{i,j})\|_{\infty}+(2\log d_{A}+1)X\|t\|_{2}\|s\|_{2})$
$\leq e^{2}(2\|(a_{i,j})\|_{\infty}+(2\log d_{A}+1)X)$
and the assertion follows by Corollary 4.3. ∎
5. Proof of Theorem 1.6
By Remark 1.7 we may assume that $a_{i,i}=0$ for all $i$ and $n\geq d_{A}\geq 3$.
For $k=1,2,\ldots$ and $t,s\in B_{2}^{n}$ we define
$I_{k}(s):=\{i\colon\ d_{A}^{-k/40}<|s_{i}|\leq d_{A}^{(1-k)/40}\},\quad J_{l}(t):=\{j\colon\ d_{A}^{-l/40}<|t_{j}|\leq d_{A}^{(1-l)/40}\}.$
Then
(5.1) $\sum_{k\geq 1}d_{A}^{-k/20}|I_{k}(s)|\leq\|s\|_{2}^{2},\quad\sum_{l\geq 1}d_{A}^{-l/20}|J_{l}(t)|\leq\|t\|_{2}^{2}$
and
${\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j\leq n}\|={\mathbb{E}}\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k,l\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$
Observe that for any $s,t\in B_{2}^{n}$,
$\Bigl{|}\sum_{k\geq 1}\sum_{l\geq k+41}\sum_{i\in I_{k}(s)}\sum_{j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{|}$ $\leq\sum_{k\geq 1}\sum_{i\in I_{k}(s)}|s_{i}|\sum_{l\geq k+41}\sum_{j\in J_{l}(t)}|a_{i,j}||t_{j}|$
$\leq\sum_{k\geq 1}\sum_{i\in I_{k}(s)}|s_{i}|d_{A}^{-(k+40)/40}\sum_{j}|a_{i,j}|$
$\leq\|(a_{i,j})\|_{\infty}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}s_{i}^{2}\leq\|(a_{i,j})\|_{\infty}.$
Similarily,
$\Bigl{|}\sum_{l\geq 1}\sum_{k\geq l+41}\sum_{i\in I_{k}(s)}\sum_{j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{|}\leq\|(a_{i,j})\|_{\infty}.$
Hence it is enough to estimate
$\sum_{r=-40}^{40}{\mathbb{E}}\sup_{\|s\|_{2}\leq}\sup_{\|t\|_{2}\leq 1}\sum_{\begin{subarray}{c}k,l\geq 1\\ l-k=r\end{subarray}}\sum_{i\in I_{k}(s)}\sum_{j\in J_{l}(t)}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$
By symmetry it is enough to bound only the terms with $r\geq 0$. Let $X$ be defined by (4.2).
Then for a fixed $r\geq 0$ and $\alpha>0$,
$\sup_{\|s\|_{2}\leq 1}$ $\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\{|J_{k+r}(t)|\geq\alpha|I_{k}(s)|\}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\leq\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\max_{\eta_{i},\eta_{j}^{\prime}=\pm 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}d_{A}^{(2-2k-r)/40}\mathbbm{1}_{\{|J_{k+r}(t)|\geq\alpha|I_{k}(s)|\}}a_{i,j}\varepsilon_{i,j}\eta_{i}\eta_{j}^{\prime}$
$\leq X\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}d_{A}^{(2-2k-r)/40}\sqrt{|I_{k}(s)||J_{k+r}(t)|}\mathbbm{1}_{\{|J_{k+r}(t)|\geq\alpha|I_{k}(s)|\}}$
$\leq\alpha^{-1/2}X\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}d_{A}^{(2-2k-r)/40}|J_{k+r}(t)|\leq\alpha^{-1/2}d_{A}^{(r+2)/40}X,$
where the last inequality follows by (5.1).
Hence Corollary 4.3 yields
${\mathbb{E}}\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\{|J_{k+r}(t)|\geq d_{A}^{(2r+5)/40}|I_{k}(s)|\}}$ $a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\log n).$
In a similar way we show that
$\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}$ $\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\{|J_{k+r}(t)|\leq\alpha|I_{k}(s)|\}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\leq\alpha^{1/2}X\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}d_{A}^{(2-2k-r)/40}|I_{k}(s)|\leq\alpha^{1/2}d_{A}^{(2-r)/40}X$
and
${\mathbb{E}}\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\{|J_{k+r}(t)|\leq d_{A}^{(2r-5)/40}|I_{k}(s)|\}}$ $a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\log n).$
Hence it is enough to bound for $r=0,1,\ldots,40$,
${\mathbb{E}}\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\left\{d_{A}^{(2r-5)/40}<|J_{k+r}(t)|/|I_{k}(s)|<d_{A}^{(2r+5)/40}\right\}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$
Recall definition (4.1) of sets $I^{\prime}$ and $I^{\prime\prime}$. Let us fix $0\leq r\leq 40$ and $k,t,s$ such that $d_{A}^{(2r-5)/40}<|J_{k+r}(t)|/|I_{k}(s)|<d_{A}^{(2r+5)/40}$. Let $|I_{k}(s)|=m$. Let us consider the following greedy algorithm with output being a subset $\{i_{1},\ldots,i_{M}\}$ of $I_{k}(s)$ of size $M\leq m$
  • In the first step we pick a vertex $i_{1}\in I_{k}(s)$ with maximal number of neighbours in $J_{k+r}(t)$.
  • Once we have $\{i_{1},\ldots,i_{N}\}$ and $N<M$, we pick $i_{N+1}\in I_{k}(s)\setminus\{i_{1},\ldots,i_{N}\}$ with maximal number of neighbours in $J_{k+r}(t)\setminus\{i_{1},\ldots,i_{N}\}^{\prime}$.
If $l_{N}$ is the number of neighbours of $i_{N}$ in $J_{k+r}(t)\setminus\{i_{1},\ldots,i_{N-1}\}^{\prime}$, then $l_{1}\geq l_{2}\geq\ldots\geq l_{M}$, so $Ml_{M}\leq|J_{k+r}(t)|$. Choose $M\leq m$ to be the maximal integer so that $l_{M}\geq d_{A}^{(r+18)/40}$ and set $I:=\{i_{1},\ldots,i_{M}\}$. This way we construct a subset $I\subset I_{k}(s)$ with cardinality $|I|\leq d_{A}^{-(r+18)/40}|J_{k+r}(t)|\leq d_{A}^{(r-13)/40}m$ such that for every $i\in I_{k}(s)\setminus I$, $|\{j\in J_{k+r}(t)\setminus I^{\prime}\colon i\sim_{A}j\}|<d_{A}^{(r+18)/40}$. Note that if $(i,j)\in E_{A}$ and $(i,j)\in(I_{k}(s)\times J_{k+r}(t))\setminus(I^{\prime\prime}\times I^{\prime})$, then $j\in J_{k+r}(t)\setminus I^{\prime}$. Therefore,
$\sum_{(i,j)\in I_{k}(s)\times J_{k+r}(t))\setminus(I^{\prime\prime}\times I^{\prime})}|a_{i,j}||s_{i}t_{j}|$ $\leq\|(a_{i,j})\|_{\infty}\sum_{i\in I_{k}(s)}|s_{i}|d_{A}^{(r+18)/40}d_{A}^{(1-k-r)/40}$
(5.2) $\leq\|(a_{i,j})\|_{\infty}d_{A}^{19/40}\sum_{i\in I_{k}(s)}s_{i}^{2}.$
Let
$s^{\prime}=(s^{\prime}_{i})_{i\in I^{\prime\prime}\cap I_{k}(s)},\quad t^{\prime}=(t^{\prime}_{j})_{j\in I^{\prime}\cap J_{k+r}(s)},$
where
$s^{\prime}_{i}:=\frac{s_{i}}{\|(s_{i})_{i\in I_{k}(s)}\|_{2}},\quad t^{\prime}_{j}:=\frac{t_{j}}{\|(t_{j})_{j\in J_{k+r}(t)}\|_{2}}.$
Then
$\|s^{\prime}\|_{2}\leq 1,\quad\|s^{\prime}\|_{\infty}\leq d_{A}^{1/40}|I_{k}(s)|^{-1/2}=d_{A}^{1/40}m^{-1/2},$
$\|t^{\prime}\|_{2}\leq 1,\quad\|t^{\prime}\|_{\infty}\leq d_{A}^{1/40}|J_{k+r}(t)|^{-1/2}\leq d_{A}^{(7-2r)/80}m^{-1/2}.$
Hence,
(5.3) $\sum_{(i,j)\in(I_{k}(s)\times J_{k+r}(t))\cap(I^{\prime\prime}\times I^{\prime})}\varepsilon_{i,j}a_{i,j}s_{i}t_{j}\leq Y_{m,r}\|(s_{i})_{i\in I_{k}(s)}\|_{2}\|(t_{j})_{j\in J_{k+r}(t)}\|_{2},$
where
$Y_{m,r}:=\max_{|I|\leq d_{A}^{(r-13)/40}m}\sup_{s\in B_{2}^{n}\cap d_{A}^{1/40}m^{-1/2}B_{\infty}^{n}}\ \sup_{t\in B_{2}^{n}\cap d_{A}^{(7-2r)/80}m^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$
Define $Y_{r}:=\max_{1\leq m\leq n}Y_{m,r}$. The main advantage of introducing the variables $Y_{r}$ is that in their definition, the suprema are taken over delocalized vectors from the sphere of dimension $|I|$ with all coordinates bounded essentially by $d_{A}^{C}|I|^{-1/2}$, and Proposition 3.3 can be used to estimate these quantities.
Estimates (5.2) and (5.3) yield
${\mathbb{E}}$ $\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\sum_{k\geq 1}\sum_{i\in I_{k}(s)}\sum_{j\in J_{k+r}(t)}\mathbbm{1}_{\Big{\{}d_{A}^{(2r-5)/40}<|J_{k+r}(t)|/|I_{k}(s)|<d_{A}^{(2r+5)/40}\Big{\}}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\leq\sup_{\|s\|_{2}\leq 1}\sup_{\|t\|_{2}\leq 1}\Biggl{(}\sum_{k\geq 1}d_{A}^{19/40}\|(a_{i,j})\|_{\infty}\sum_{i\in I_{k}(s)}s_{i}^{2}+{\mathbb{E}}Y_{r}\sum_{k\geq 1}\|(s_{i})_{i\in I_{k}(s)}\|_{2}\|(t_{j})_{j\in J_{k+r}(t)}\|_{2}\Biggr{)}$
$\leq d_{A}^{19/40}\|(a_{i,j})\|_{\infty}+{\mathbb{E}}Y_{r}\sup_{\|s\|_{2}\leq 1}\Bigl{(}\sum_{k\geq 1}\|(s_{i})_{i\in I_{k}(s)}\|_{2}^{2}\Bigr{)}^{1/2}\sup_{\|t\|_{2}\leq 1}\Bigl{(}\sum_{k\geq 1}\|(t_{j})_{j\in J_{k+r}(t)}\|_{2}^{2}\Bigr{)}^{1/2}$
$\leq d_{A}^{19/40}\|(a_{i,j})\|_{\infty}+{\mathbb{E}}Y_{r}.$
Therefore, to establish Theorem 1.6 it is enough to prove the following lemma.
Lemma 5.1.
For every $0\leq r\leq 40$,
${\mathbb{E}}Y_{r}={\mathbb{E}}\max_{1\leq m\leq n}Y_{m,r}\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\log n)+\mathrm{Log}(d_{A})\|(a_{i,j})\|_{\infty}.$
First we show a connected counterpart to Lemma 5.1.
Lemma 5.2.
We have
${\mathbb{E}}\max_{1\leq k\leq n}\max_{I\in\mathcal{I}_{4}(k)}\ \sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}\$$\sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\log n)+\mathrm{Log}(d_{A})\|(a_{i,j})\|_{\infty}.$
Proof.
Let us first fix $k$ and $I\in\mathcal{I}_{4}(k)$. Then $|I^{\prime}|\leq d_{A}k$ and $|I^{\prime\prime}|\leq d_{A}^{2}k$. Proposition 3.3, applied with $(a_{i,j})=(a_{i,j})_{i\in I^{\prime\prime},j\in I^{\prime}}$, $n=|I^{\prime\prime}|$, $m=|I^{\prime}|$, and $b=d_{A}^{3/8}k^{-1/2}$ yields
${\mathbb{E}}\sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}\ \sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}$ $a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+\mathrm{Log}(d_{A})\|(a_{i,j})\|_{\infty}.$
By Lemma 3.2, $|\mathcal{I}_{4}(k)|\leq n(4d_{A}^{4})^{k}\leq\max\{n^{2},d_{A}^{12k}\}$ (recall that we assume that $d_{A}\geq 3$). We have
$\sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}\ \sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\Bigl{\|}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{\|}_{\mathrm{Log}(|\mathcal{I}_{4}(k)|)}$
$\leq\sup_{s,t\in B_{2}^{n}}\Bigl{\|}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{\|}_{2\mathrm{Log}(n)}+\sup_{s\in B_{2}^{n}}\sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\Bigl{\|}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{\|}_{12k\mathrm{Log}(d_{A})}$
$\leq R_{A}(2\mathrm{Log}n)+\sup_{s\in B_{2}^{n}}\sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\sqrt{12k\mathrm{Log}(d_{A})}\Bigl{(}\sum_{i,j}a_{i,j}^{2}s_{i}^{2}t_{j}^{2}\Bigr{)}^{1/2}$
$\lesssim R_{A}(\mathrm{Log}\,n)+\sqrt{k\mathrm{Log}(d_{A})}\max_{i}d_{A}^{-1/16}k^{-1/2}\Bigl{(}\sum_{j}a_{i,j}^{2}\Bigr{)}^{1/2}$
$\leq R_{A}(\mathrm{Log}\,n)+\sqrt{\mathrm{Log}(d_{A})}d_{A}^{-1/16}\max_{i}\|(a_{i,j})_{j}\|_{2}\lesssim R_{A}(\log n)+\max_{i}\|(a_{i,j})_{j}\|_{2}.$
Hence, by Proposition 3.1,
${\mathbb{E}}\max_{I\in\mathcal{I}_{4}(k)}\sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}\$$\sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$
$\lesssim\max_{i}\|(a_{i,j})_{j}\|_{2}+R_{A}(\log n)+\mathrm{Log}(d_{A})\|(a_{i,j})\|_{\infty}.$
Applying again Proposition 3.1 and observing that
$\max_{1\leq k\leq n}\sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}\ \sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \Bigl{\|}\sum_{i,j}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}\Bigr{\|}_{\mathrm{Log}(n)}\leq R_{A}(\log n)$
we get the assertion. ∎
Proof of Lemma 5.1.
Let
$Z_{k}:=\max_{I\in\mathcal{I}_{4}(k)}\sup_{s\in B_{2}^{n}\cap d_{A}^{3/8}k^{-1/2}B_{\infty}^{n}}\ \sup_{t\in B_{2}^{n}\cap d_{A}^{-1/16}k^{-1/2}B_{\infty}^{n}}\ \sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$
Let us fix $0\leq r\leq 40$, $I\subset V$ such that $|I|\leq d_{A}^{(r-13)/40}m$, $s\in B_{2}^{n}\cap d_{A}^{1/40}m^{-1/2}B_{\infty}^{n}$ and $t\in B_{2}^{n}\cap d_{A}^{(7-2r)/80}m^{-1/2}B_{\infty}^{n}$. Let $I_{1},\ldots,I_{l}$ be 4-connected components of $I$. Then $\{I_{1}^{\prime},\ldots,I_{l}^{\prime}\}$ is a partition of $I^{\prime}$, $\{I_{1}^{\prime\prime},\ldots,I_{l}^{\prime\prime}\}$ is a partition of $I^{\prime\prime}$ and
(5.4) $\sum_{i\in I^{\prime\prime},j\in I^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}=\sum_{u=1}^{l}\sum_{i\in I_{u}^{\prime\prime},j\in I_{u}^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}.$
Define
$U_{1}$ $:=\bigl{\{}1\leq u\leq l\colon\ \|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}\geq d_{A}^{(12-r)/80}m^{-1/2}\sqrt{|I_{u}|}\bigr{\}},$
$U_{2}$ $:=\bigl{\{}1\leq u\leq l\colon\ \|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}\geq d_{A}^{(12-r)/80}m^{-1/2}\sqrt{|I_{u}|}\bigr{\}}.$
For $u\in U_{1}\cap U_{2}$ define vectors
$\tilde{s}(u):=\frac{(s_{i})_{i\in I_{u}^{\prime\prime}}}{\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}},\quad\tilde{t}(u):=\frac{(t_{j})_{j\in I_{u}^{\prime}}}{\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}}.$
Then $\|\tilde{s}(u)\|_{2}=\|\tilde{t}(u)\|_{2}=1$,
$\|\tilde{s}(u)\|_{\infty}$ $\leq d_{A}^{(r-12)/80}m^{1/2}|I_{u}|^{-1/2}\|s\|_{\infty}\leq d_{A}^{(r-10)/80}|I_{u}|^{-1/2}\leq d_{A}^{3/8}|I_{u}|^{-1/2},$
$\|\tilde{t}(u)\|_{\infty}$ $\leq d_{A}^{(r-12)/80}m^{1/2}|I_{u}|^{-1/2}\|t\|_{\infty}\leq d_{A}^{-(r+5)/80}|I_{u}|^{-1/2}\leq d_{A}^{-1/16}|I_{u}|^{-1/2}.$
Hence
$\sum_{u\in U_{1}\cap U_{2}}\sum_{i\in I_{u}^{\prime\prime},j\in I_{u}^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$ $\leq\sum_{u\in U_{1}\cap U_{2}}Z_{|I_{u}|}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}$
$\leq\max_{k}Z_{k}\Bigl{(}\sum_{u\leq l}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}^{2}\Bigr{)}^{1/2}\Bigl{(}\sum_{u\leq l}\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}^{2}\Bigr{)}^{1/2}$
(5.5) $\leq\max_{k}Z_{k}.$
Observe that
$\sum_{u\notin U_{1}}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}^{2}\leq\sum_{u}d_{A}^{(12-r)/40}m^{-1}|I_{u}|=d_{A}^{(12-r)/40}m^{-1}|I|\leq d_{A}^{-1/40}$
and by the same token
$\sum_{u\notin U_{2}}\|(t_{j})_{i\in I_{u}^{\prime}}\|_{2}^{2}\leq d_{A}^{-1/40}.$
Hence
$\sum_{u\notin U_{1}}\sum_{i\in I_{u}^{\prime\prime},j\in I_{u}^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$ $\leq\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|\sum_{u\notin U_{1}}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}$
$\leq\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|\Bigl{(}\sum_{u\notin U_{1}}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}^{2}\Bigr{)}^{1/2}\Bigl{(}\sum_{u\leq l}\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}^{2}\Bigr{)}^{1/2}$
(5.6) $\leq d_{A}^{-1/80}\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|$
and
$\sum_{u\in U_{1}\setminus U_{2}}\sum_{i\in I_{u}^{\prime\prime},j\in I_{u}^{\prime}}a_{i,j}\varepsilon_{i,j}s_{i}t_{j}$ $\leq\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|\sum_{u\notin U_{2}}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}$
$\leq\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|\Bigl{(}\sum_{u\leq l}\|(s_{i})_{i\in I_{u}^{\prime\prime}}\|_{2}^{2}\Bigr{)}^{1/2}\Bigl{(}\sum_{u\notin U_{2}}\|(t_{j})_{j\in I_{u}^{\prime}}\|_{2}^{2}\Bigr{)}^{1/2}$
(5.7) $\leq d_{A}^{-1/80}\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|.$
Bounds (5.4)-(5.7) yield
${\mathbb{E}}\max_{m}Y_{m,r}\leq{\mathbb{E}}\max_{k}Z_{k}+2d_{A}^{-1/80}{\mathbb{E}}\|(a_{i,j}\varepsilon_{i,j})_{i,j}\|$
and the assertion follows by Lemma 5.2 and Proposition 4.4. ∎
Acknowledgements. Part of this work was carried out while the author was visiting the Hausdorff Research Institute for Mathematics, Univerity of Bonn. The hospitality of HIM and of the organizers of the program Synergies between modern probability, geometric analysis and stochastic geometry is gratefully acknowledged. The author would like to thank Marta Strzelecka for careful readings of various versions of the manuscript and numerous valuable comments, Ramon van Handel for pointing out the reference [7] and an anonymous referee for many thoughtful comments and for suggesting most of the examples presented in the paper.